### 07 Feb 2024 v1

# Optimistic Thompson Sampling for No-Regret Learning in Unknown Games

Abstract
Many real-world problems involving multiple decision-makers can be modeled as an unknown game characterized by partial observations. Addressing the challenges posed by partial information and the curse of multi-agency, we developed Thompson sampling-type algorithms, leveraging information about opponent’s action and reward structures. Our approach significantly reduces experimental budgets, achieving a more than tenfold reduction compared to baseline algorithms in practical applications like traffic routing and radar sensing. We demonstrate that, under certain assumptions about the reward structure, the regret bound exhibits merely a logarithmic dependence on the total action space size, effectively mitigating the curse of multi-agency. Additionally, this research introduces the Optimism-then-NoRegret framework, a novel contribution that integrates both our proposed methodologies and existing algorithms in the field.
Machine Learning, ICML
Figure 1: Samples required for different algorithms to reach the same average regret $(10^{-1})$ under a $50\times 50$ matrix with the self-play setting. (UCB-RM and OTS type are proposed algorithms. Moreover, IWE-Hedge corresponds to the Exp3 algorithm, and IWE-RM represents regret matching with bandit feedback. The UCB-Hedge algorithm was introduced by (Sessa et al., 2019).)
1 Introduction
Many real-world problems in economics (Fainmesser, 2012), sociology (Skyrms & Pemantle, 2009), transportation (Leblanc, 1975), politics (Ordeshook et al., 1986), signal processing (Song et al., 2011), and other fields (Fudenberg & Tirole, 1991) can be described as unknown games, where each player only observes their opponents’ actions and the noisy rewards associated with their own selected actions (referred to as bandit feedback). The goal of each player is to maximize their individual reward, and the only way to achieve this is to repeatedly play and learn the game structure from the corresponding observed rewards. The challenge in unknown games is how to efficiently learn from bandit feedback. Celebrated no-regret learning algorithms, such as Hedge (Freund & Schapire, 1997) and regret-matching (Hart & Mas-Colell, 2000), ensure sublinear regret guarantees under the full information setting, where the rewards of all actions at each round are observable. However, these algorithms cannot handle problems with bandit feedback, and faces the curse of multi-agency. This paper focuses on how to effectively utilize information on opponents’ actions and reward structures, which can substantially reduce budgets, as the example illustrated in Fig.1.
Table 1: Regret bounds comparison.
Feedback Full Bandit Bandit + Actions Bandit + Actions
Imagined Reward IWE UCB OTS [Ours]
No-Regret Update Hedge $\mathcal{O}\big{(}\sqrt{T\log\mathcal{A}}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\mathcal{A}\log\mathcal{A}}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\log\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\log\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$
RM $\mathcal{O}\big{(}\sqrt{T\mathcal{A}}\big{)}$ $\mathcal{O}\big{(}T^{2/3}\mathcal{A}^{2/3}\big{)}$ [Ours] $\mathcal{O}\big{(}\sqrt{T\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$ [Ours] $\mathcal{O}\big{(}\sqrt{T\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$
Table 2: $\mathcal{A}$ and $\mathcal{B}$ are the sizes of action sets, $\gamma_{T}$ is the maximum information gain (Definition 4.3) and $\beta=\log(\mathcal{A}T)$.
1.1 Related Works
Adversarial bandit
In the full-information setting, multiplicative-weights (MW) algorithms such as Hedge (Freund & Schapire, 1997) can achieve optimal regret for adversarial bandit problems. Full information feedback requires perfect game knowledge and is unrealistic in many applications. In the more challenging bandit setting, Exp3 (Auer et al., 2002b) is a classical algorithm that utilizes the importance-weighted estimator for reward vector construction.
Game
[…]
Previous literature ignores the fact that the agents are playing repeated games in which the reward structure could be exploited. Furthermore, to the application of our interest, the opponent’s actions can be observed by the agent. This scenario has received relatively little attention despite its numerous applications (O’Donoghue et al., 2021). Some papers consider a similar setting as ours. (O’Donoghue et al., 2021) seeks to compare the received reward to the Nash value and proposes variants of UCB and K-learning that converge to the Nash value. Our goal is to exploit the weakness of opponent and compete with hindsight’s best actions. With this performance metric, we seek to exploit the opponent’s strategy instead of only achieving Nash value. Similar to our setting, (Sessa et al., 2019) also uses adversarial regret and focuses on the single-player viewpoint. It utilizes Gaussian Process to exploit the correlations among different game outcomes and obtains a kernel-dependent regret bound with the factor $\gamma_{T}$ by an upper confidence bound (UCB) type algorithm. Thompson Sampling (TS) and other randomized exploration variants (Russo et al., 2018; Vaswani et al., 2020) are a strong counterpart of UCB-type algorithm in the reward structure-aware bandit literature. Moreover, (Zhang, 2022; Agarwal & Zhang, 2022) implements the optimism in Thompson sampling by introducing a optimism prior, which favors high reward models more aggressively than the standard Thompson sampling. However, this approach for optimistic TS is not computationally tractable. Another line of work is to tackle computation tractability issue of Thompson sampling (Li et al., 2022) for large-scale complex environment, but it remains the single-agent setups. However, no evidence shows these randomized exploration methods can work in our unknown game setting.
This work tries to understand how to efficiently learn to play unknown games. The main contributions are highlighted as follows:
  • Effective algorithm: We first demonstrate that vallina TS indeed fails in a specific class of unknown games. To overcome this, we introduce an optimistic variant of TS (referred as OTS for short) combined with appropriate full information adversarial bandit algorithms. We show that OTS can fix the divergence issue with the help of anti-concentration behavior of Gaussian distribution. Moreover, it significantly reduces experimental budgets in two practical problems. It decreases the budgets more than tenfold [^1] In Figure 5, OTS-Hedge achieves more than tenfold improvement comparing with IWE-hedge for reaching regret $10^{-2}$. compared to baseline algorithms in a two player radar sensing problem, and effectively mitigates the curse of multi-agency in traffic routing problem up to hundreds of decision-makers [^2] In Figure 6, OTS type and UCB type algorithms can effective handle hundreds of decision-makers..
  • Theoretical advantages: A general information-theoretic regret bound is provided and sublinear regret bounds of all proposed algorithms are established accordingly (Table 2). Our analysis highlights that using (i) the information of opponent’s actions and (ii) the underlying reward structure can help resolve the curse of multi-agency. For structured reward functions, our algorithms achieve regret bounds that depend logarithmically on the size of the action space. In contrast, algorithms that rely only on partial information suffer from curse of multi-agency.
  • Framework: An Optimism-then-NoRegret (OTN) learning framework for unknown games is also introduced. This framework encompasses various vanilla game algorithms as special cases, and several efficient algorithms can be developed under this framework, including upper confidence bound (UCB) based algorithms and TS-based algorithms. Notably, the proposed OTS-RM algorithm from this framework achieves the best performance in all experiments.
This paper is organized as follows. Section 2 introduces the fundamental protocol and notations used in the repeated bandit game. Section 3 describes the proposed OTS type and UCB-RM algorithms, along with the corresponding OTN framework. Section 4 provides a regret analysis, while Section 5 details the experimental results. Section 6 concludes the paper.
2 Repeated Bandit Game
To simplify the presentation, we consider a two-player game scenario involving Alice and Bob. However, our results can be extended straightforwardly to multiplayer games by treating all other players as an abstract player.
Protocol.
Consider a repeated game between Alice and her opponent Bob, where the action index sets for Alice and Bob are denoted by $\mathcal{A}={1,\ldots,\left|\mathcal{A}\right|}$ and $\mathcal{B}={1,\ldots,\left|\mathcal{B}\right|}$, respectively. [^3] We use the shorthand $\mathcal{A}$ to denote the cardinality $\left|\mathcal{A}\right|$. At each time $t=0,1,\ldots$, Alice selects an action $A_{t}\in\mathcal{A}$, and Bob simultaneously selects an action $B_{t}\in\mathcal{B}$. The mean reward for each action pair $(a,b)$ is $f_{\theta}(a,b)$, where $\theta\in\Theta$ is a model parameter unknown to the players. In the bandit feedback case, Alice only observes a noisy version of the mean reward associated with the selected action pair $(A_{t},B_{t})$
$Y_{t+1,A_{t},B_{t}}=f_{\theta}(A_{t},B_{t})+\eta_{t},$
where $\eta_{t}$ is an iid noise sequence. Under the full information setting, Alice can observe the whole mean rewards $\textbf{r}_{t}=(f_{\theta}(a,B_{t}))_{a\in\mathcal{A}}$ (details are presented in Appendix B). Alice’s experience up to time $t$ is encoded by the history $H_{t}=\left(A_{0},B_{0},Y_{1,A_{0},B_{0}},\ldots,A_{t-1},B_{t-1},Y_{t,A_{t-1},B_{t-1}}\right)$.
Algorithm.
An algorithm $\pi^{\operatorname{alg}}=(\pi_{t})_{t\in\mathbb{N}}$ employed by Alice is a sequence of deterministic functions, where each $\pi_{t}(H_{t})$ specifies a probability distribution over the action set $\mathcal{A}$ based on the history $H_{t}$. Alice’s action $A_{t}$ is sampled from the distribution $\pi_{t}$, i.e., $\mathbb{P}(A_{t}\in\cdot\>|\>\pi_{t})=\mathbb{P}(A_{t}\in\cdot\>|\>H_{t})=\pi_{t}(\cdot)$.
The above description of the bandit game encompasses various game forms based on the structure of the mean function $f_{\theta}(a,b)$. Several representative game forms such as the matrix game, linear game, and kernelized game are summarized in Table 4 (see Appendix B).
Reward and performance metric.
[…]
$\sum_{t=0}^{T-1}\mathbb{E}\left[R_{t+1,A^{*},B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right],$ (1)
where the expectation is taken over the randomness in the actions $A_{t}$ and the rewards $R_{t+1,A_{t},B_{t}}$. However, this adversarial regret $\Re(T,\pi^{\operatorname{alg}},B_{0:T},\theta)$ is not a suitable regret metric under our game setting since it depends on the specific action sequence $B_{0:T}$ played by Bob. In this paper, we adopt the worse-case regret as the metric. An algorithm $\pi_{\operatorname{alg}}$ is considered No-Regret for Alice if, for any $B_{0:T}$, Alice suffers only sublinear regret, i.e.,
*$\Re^{*}(T,\pi^{\operatorname{alg}})=\sup_{B_{0:T}\in\mathcal{B}^{T}}\Re(T,\pi^{\operatorname{alg}},B_{0:T})=o(T),$
where $\theta$ is omitted for notation simplicity.
3 Optimism-then-NoRegret Learning
3.1 Review of Full Information Feedback
We will start by providing a brief overview of the full information feedback setting, in which Alice can observe the mean rewards $r_{t}(a)=f_{\theta}(a,B_{t})$ for all actions $a\in\mathcal{A}$ (details are presented in Appendix B). At time $t$, Alice picks action $A_{t}\sim P_{X_{t}}$ where $P_{X}(i)={X_{i}}/\sum_{i\in\mathcal{A}}X_{i}$. Full-information adversarial bandit algorithms, e.g., Hedge (Freund & Schapire, 1997) and Regret Matching (RM) (Hart & Mas-Colell, 2000), can be used to update $X_{t}$ to $X_{t+1}=g_{t}(X_{t},r_{t})$,
$\begin{cases}\textrm{Hedge:}\ &g_{t,a}(X_{t},r_{t})=X_{t,a}\exp\left(\eta_{t}r_{t}(a)\right),\\ \textrm{RM:}\ &g_{t,a}(X_{t},r_{t})=\max\left(0,\sum\limits_{s=0}^{t}r_{t}(a)-r_{t}(A_{s})\right),\end{cases}$
where $g_{t}(\cdot):\mathbb{R}_{+}^{\mathcal{A}}\times\mathbb{R}_{+}^{\mathcal{A}}\mapsto\mathbb{R}_{+}^{\mathcal{A}}$. The full information adversarial regret of an algorithm $\operatorname{adv}$ for a reward sequence $(r_{t})_{t}$ is defined as
$\Re_{\operatorname{full}}(T,{\operatorname{adv}},(r_{t})_{t})=\max_{a\in\mathcal{A}}\mathbb{E}\left[\sum_{t=0}^{T-1}r_{t}(a)-r_{t}(A_{t})\right].$ (2)
The worst-case regret is defined as
$\Re_{\operatorname{full}}(T,{\operatorname{adv}})$ $=\max_{(r_{t})_{t}}\Re_{\operatorname{full}}(T,{\operatorname{adv}},(r_{t})_{t}).$ (3)
Since $r_{t}(a)=f_{\theta}(a,B_{t})$, the adversarial regret in Equation 1 can be reformulated as the full-information adversarial regret. For Hedge and RM (Freund & Schapire, 1997), their worst regret regrets can be bounded as $\Re_{\operatorname{full}}(T,{\operatorname{Hedge}})=\mathcal{O}(\sqrt{T\log\mathcal{A}})$ and $\Re_{\operatorname{full}}(T,{\operatorname{RM}})=\mathcal{O}(\sqrt{T\mathcal{A}})$.
3.2 Bandit Feedback
In the bandit feedback setting, Alice can only observe . We propose a framework that combines Optimism algorithm for the stochastic bandits and Noregret algorithm for the full information adversarial bandits. First, we construct a sequence of surrogate full information feedback $\tilde{R}_{t}\in\mathbb{R}^{\mathcal{A}}$ in an optimistic sense, which we refer to as the imagined reward. In particular, we can use an optimism algorithm $E$ for stochastic bandit
$\tilde{R}_{t+1}=E(H_{t+1},Z_{t+1})\in\mathbb{R}^{\mathcal{A}}.$ (4)
Then, we apply a noregret update rule $g_{t}$ to update the sampling distribution with $\tilde{R}_{t+1}$
$X_{t+1}=g_{t}(X_{t},\tilde{R}_{t+1}).$ (5)
This procedure is described in Algorithm 1, termed as Optimism-then-NoRegret (OTN).
Algorithm 1 Optimism-then-NoRegret
1:  Initialize $X_{1}$
2:  for round $t=0,1,\ldots,T-1$  do
3:     Sample action $A_{t}\sim P_{X_{t}}$
4:     Observe opponent’s action $B_{t}$ and noisy bandit feedback $R_{t+1,A_{t},B_{t}}$
5:     Update history: $H_{t+1}=(H_{t},A_{t},B_{t},R_{t+1,A_{t},B_{t}})$
6:     Construct reward vector $\tilde{R}_{t+1}$ with an optimism algorithm $E$
7:     Noregret update
8:  end for
The essential part in Algorithm 1 is the construction of the imagined reward vector $\tilde{R}_{t+1}$.
To motivate, we provide a key regret decomposition
Proposition 3.1 (Regret decomposition).
For any $a\in\mathcal{A}$, the one-step regret can be decomposed as
$\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right]=(I)+(II)+(III)$
where
$(I)$ $=\mathbb{E}\left[\tilde{R}_{t+1}(a)-\tilde{R}_{t+1}(A_{t})\mid\theta\right],$
$(II)$ $=\mathbb{E}\left[f_{\theta}({a,B_{t}})-\tilde{R}_{t+1}(a)\mid\theta\right],$
$(III)$ $=\mathbb{E}\left[\tilde{R}_{t+1}(A_{t})-f_{\theta}(A_{t},B_{t})\mid\theta\right].$
Summation of $(I)$ reduces to adversarial regret $\Re_{full}(T,\operatorname{adv},\tilde{R})$ for a bounded sequence $\tilde{R}_{t+1}$. For the term $(II)$, sufficient optimism guarantees $\sqrt{T}$-type regret. Therefore, the OTN framework can achieve sublinear regret when equipped with reasonable optimism algorithm and noregret algorithm. Next we discuss several algorithms that can be implied by Algorithm 1.
A celebrated estimation method is importance-weighted estimator (IWE) (Lattimore & Szepesvári, 2020), which combined with Hedge becomes the well known Exp3 algorithm (Auer et al., 2002b) (referred as IWE-Hedge in this paper). Additionally, IWE can also be combined with RM, resulting in IWE-RM for the bandit game. Details of IWE-Hedge and IWE-RM are presented in Appendix D, and their adversarial regrets are provided below.
Proposition 3.2 (Regrets of IWE-Hedge (Auer et al., 2002b)).
Consider play bandit game with IWE-Hedge or IWE-RM, then *$\Re^{*}(T,\text{IWE-Hedge})=\mathcal{O}(\sqrt{T\mathcal{A}\log\mathcal{A}}).$
We also extend the regret matching (RM) (Hart & Mas-Colell, 2000) to our setting under the OTN framework, resulting IWE-RM algorithm (detailed proof can be found in Section D).
Theorem 3.3 (Regret of IWE-RM).
Consider play bandit game with IWE-RM, then *$\Re^{*}(T,\text{IWE-RM})=\mathcal{O}(T^{2/3}\mathcal{A}^{2/3}).$
Remark 3.4.
IWE-RM is our by-product result which is not known in the literature.
However, IWE only requires the bandit reward feedback and does not explicitly utilize the information of opponents’ action. An important idea is that, with the information of opponent’s actions and the knowledge on reward structure $f_{\theta}$, we can estimate $f_{\theta}(a,b)$ for all $(a,b)\in\mathcal{A}\times\mathcal{B}$ by $\tilde{f}_{t+1}(a,b)$ at each time $t$, and construct imagined reward by $\tilde{R}_{t+1}(a)=\min(\tilde{f}_{t+1}(a,B_{t}),1)$. By taking Gaussian distribution as the prior, the mean and variance estimation of $f_{\theta}(a,b)$ up to time $t$ can be recursively updated, denoted as $\mu_{t}(a,b)$ and $\sigma_{t}^{2}(a,b)$. Detailed updating formulas for games in Table 4 are presented in Appendix C. The remaining issue is how to balance exploration and exploitation based on mean and variance estimation.
In stochastic bandit, UCB (Auer et al., 2002a) and TS (Thompson, 1933) are two widely used techniques to achieve optimism. Tailoring to the bandit game, the corresponding constructions using UCB and TS are as follows with $\beta_{t}=\sqrt{2\log{\mathcal{A}\sqrt{t}}}$:
$\begin{cases}\textrm{UCB:}&\tilde{f}_{t+1}(a,B_{t})\>|\>H_{t+1}=\mu_{t}(a,B_{t})+\beta_{t}\sigma_{t}(a,B_{t}),\\ &\tilde{R}_{t+1}(a)=\tilde{f}_{t+1}(a,B_{t})\wedge 1,\forall a\in\mathcal{A}.\\ \textrm{TS:}&\tilde{f}_{t+1}(a,B_{t})\>|\>H_{t+1}\sim N(\mu_{t}(a,B_{t}),\sigma_{t}(a,B_{t})),\\ &\tilde{R}_{t+1}(a)=\tilde{f}_{t+1}(a,B_{t})\wedge 1,\forall a\in\mathcal{A}.\end{cases}$
The effectiveness of UCB combined with Hedge has been studied in (Sessa et al., 2019). Within the OTN framework, UCB can also be combined with RM, resulting in a sublinear regret based on our analysis (see details in Section 4). The regrets of UCB-Hedge and UCB-RM are provided below:
Theorem 3.5 (Regrets of UCB-Hedge and UCB-RM).
Consider playing the bandit feedback game with UCB-Hedge or UCB-RM, then *$\Re^{*}(T,\text{UCB-Hedge})=\mathcal{O}(\sqrt{T\log\mathcal{A}}+\sqrt{\gamma_{T}\beta T}),\ \Re^{*}(T,\text{UCB-RM})=\mathcal{O}(\sqrt{T\mathcal{A}}+\sqrt{\gamma_{T}\beta T}).$
Remark 3.6.
Our regret analysis of UCB-Hedge is under Bayesian setting while that in (Sessa et al., 2019) is under frequentist setting. Due to the setting difference, our regret bound has $\sqrt{\gamma_{T}}$ improvement.
3.3 Failure of TS
We demonstrate that TS combined with RM fails in a class of bandit games by providing a counter example.
Example 3.1 (Best Response Player).
Consider a class of matrix games with a payoff matrix $\theta$ defined as
$\theta=\begin{bmatrix}1&1-\Delta\\ 1-\Delta&1\end{bmatrix},$ (6)
where $\Delta\in(0,1)$. At time $t$, Alice plays action $A_{t}\sim P_{X_{t}}$, and Bob is a best response player who selects action $B_{t}\sim P_{Y_{t}}$ according to the best response strategy $Y_{t}=\min_{y}X_{t}^{\top}\theta y$. Alice receives the noiseless reward $R_{t+1,A_{t},B_{t}}=\theta_{A_{t},B_{t}}$.
Figure 2: Failure probability for TS-RM and OTS-RM.
Intuitively, as long as Alice chooses pure strategy, Bob will exploit Alice, resulting in $\Delta$ regret. Specifically, let $\Omega_{t}$ denote the failure event where Alice selects the 2nd row and Bob chooses the 1st column for all time steps $t^{\prime}\leqslant t$. Assuming uniform initialization for Alice, then $\Re(T)\geqslant 2\mathbb{P}(\Omega_{T})\Delta\cdot T.$ If the event $\Omega_{t}$ occurs with a constant probability for all $t\geqslant 1$, then Alice suffers linear regret.
As shown in Figure 2, TS-RM indeed fails in different setups for hyperparameter $(\Delta,\sigma_{n}^{2})$. Detailed mathematical justification can be found in Appendix E.
Proposition 3.7 (Failure of TS-RM).
Suppose Alice initializes with a uniform strategy and utilizes Regret Matching with TS estimator (TS-RM). For any $\Delta\in(0,1)$ and $\sigma_{n}>0$, there exists a universal constant $c(\Delta,\sigma_{n})>0$ such that for all rounds $t\geqslant 1$, $\mathbb{P}(\Omega_{t})\geqslant c(\Delta,\sigma_{n})$.  [^5] If $\Delta=0.1$ and $\sigma_{n}=0.1$, we have $c(\Delta,\sigma_{n})\approx 0.54$.
The proof of Proposition 3.7 can be found in Appendix E.
3.4 A Simple Fix: Optimistic Sampling
By taking the maximum of multiple functions sampled from the posterior, we can improve the probability of being optimistic for the estimator. At time $t$, after observing $H_{t+1}=(H_{t},A_{t},B_{t},R_{t+1,A_{t},B_{t}})$, we can perform optimistic sampling by first sampling $M_{t+1}$ independent normal random samples and constructing the estimator: $\forall a\in\mathcal{A}$
$\tilde{f}^{\operatorname{TS},j}_{t+1}(a,B_{t})$ $\sim N(\mu_{t}(a,B_{t}),\sigma_{t}(a,B_{t})),j\in[M_{t+1}]$
$\tilde{f}^{\operatorname{OTS}}_{t+1}(a,B_{t})$ $:=\max_{j\in[M_{t+1}]}\tilde{f}^{\operatorname{TS},j}_{t+1}(a,B_{t}),$
$\tilde{R}^{\operatorname{OTS}}_{t+1}(a)$ $:=\operatorname{clip}_{[0,1]}\left(\tilde{f}_{t+1}^{\operatorname{OTS}}(a,B_{t})\right).$
To intuitively understand why optimistic samples help, let us recall Example 3.1. Denote $\tilde{R}^{OTS}_{t+1}(1st)$ and $\tilde{R}^{OTS}_{t+1}(2nd)$ as the optimistic reward for the first and second row actions, respectively. Then the probability of $\tilde{R}^{OTS}_{t+1}(1st)>\tilde{R}^{OTS}_{t+1}(2nd)$ increases as $M_{t+1}$ grows, this implies a decay of $\mathbb{P}(\Omega_{t})$ over time $t$. Detailed mathematical justification can be found in Appendix E. Such a simple optimistic version of TS (referred to as OTS) has the following regret guarantee.
Theorem 3.8 (Regrets of OTS-Hedge and OTS-RM).
For any full information adversarial bandit $\operatorname{adv}$ working with the imagined reward sequence $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t+1},t=0,1,\ldots)$ constructed by OTS with $M_{1}=M_{2}=\ldots=M_{T}=M=\mathcal{O}(\log\mathcal{A}T)$, the combined algorithm $\pi=\pi^{\operatorname{adv-OTS}}$ enjoys the regret bound:
*$\Re^{*}(T,\pi,\theta)\leqslant\Re_{\operatorname{full}}(T,\operatorname{adv},\tilde{R}^{\operatorname{est}})+\sqrt{\log(\mathcal{A}T)I(\theta;H_{T})T}$
where $\Re_{\operatorname{full}}$ is defined in Equation 3.
Note that as long as the imagined $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t+1},t=0,1,\ldots)$ satisfies bounded property $\tilde{R}_{t}\in[0,1]^{\mathcal{A}}$, then if we take $\operatorname{adv}$ to be Hedge, $\Re_{\operatorname{full}}(T,\text{Hedge},\tilde{R}^{\operatorname{est}})=\mathcal{O}(\sqrt{T\log\mathcal{A}})$. And if taking $\operatorname{adv}$ to be RM, then $\Re_{\operatorname{full}}(T,\text{RM},\tilde{R}^{\operatorname{est}})=\mathcal{O}(\sqrt{T\mathcal{A}})$.
Remark 3.9.
The regret bound for UCB shares the same order as OTS but there is a fundamental difference, i.e., OTS actually is a stochastic upper confidence bound method. Further, from practical implementation view, OTS has some potential benefits like obtaining an (approximate) posterior sample of complex function is computationally easier than compute the UCBs of the function.
4 Regret Analysis
Our analysis road map is as follows. First, recall that Proposition 3.1 presents a general regret decomposition
$\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right]=(I)+(II)+(III)$
Summation of $(I)$ reduces to adversarial regret $\Re_{full}(T,\operatorname{adv},\tilde{R})$ for a bounded sequence $\tilde{R}_{t+1}$. For the term $(II)$, the following definition of optimism is a sufficient condition for $\sqrt{T}$-type regret. In addition, Proposition 3.1 also applies to IWE and UCB, resulting in our by-product regret bounds of IWE-RM and UCB-RM (see Table 2 and details are presented in Appendix F).
Definition 4.1 (Sufficient optimism).
We say the constructed imagined reward sequence $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t},t\in\mathbb{Z}_{++})$ is optimistic if for any action $a\in\mathcal{A}$, $\mathbb{P}(f_{\theta}(a,B_{t})\geqslant\tilde{R}_{t+1}(a)\mid\theta)\leqslant\mathcal{O}(1/\sqrt{T}).$
OTS can satisfy Definition 4.1 by selecting proper $M_{t+1}$ whereas TS cannot satisfy. UCB sequence can also satisfy Definition 4.1. The term $(III)$ can be bounded by one-step information gain $I(\theta;R_{t+1,A_{t},B_{t}}|H_{t})$ using differential entropy of Gaussian distribution. Here, we give a general regret bound for any imagined reward sequence.
Theorem 4.2.
For any full information adversarial bandit $\operatorname{adv}$ working the imagined reward $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t},t\in\mathbb{Z}_{++})$ constructed by estimation algorithm $\operatorname{est}$, if the $\tilde{R}^{\operatorname{est}}$ satisfy Definition 4.1, the regret of the combined algorithm $\pi=\pi^{\operatorname{adv-est}}$ enjoys
*$\Re^{*}(T,\pi,\theta)$ $\leqslant\Re_{\operatorname{full}}(T,\operatorname{adv})+\sqrt{\beta I(\theta;H_{T})T},$
where $\Re_{\operatorname{full}}$ is defined in Equation 3 and $\beta=\mathcal{O}(\log\mathcal{A}T)$.
Proof sketch. First, we derive a general regret decomposition for any imagined reward sequence, and bound each term separably. We bound the first term via a reduction to adversarial regret. For the second term, we introduce the generic UCB sequence and lower confidence bound (LCB) sequence and derive a general regret bound for these sequences. We bound the third term via information-theoretic quantity. Finally, combining the three bounds together yields the regret upper bound. The detailed proof is provided in Appendix F.
To characterize how much uncertainty reduction is in a particular game structure when observing new information, we define the maximum information gain.
Definition 4.3 (Maximum Information Gain).
$\gamma_{T}:=\max_{A_{0:T},B_{0:T}}I(\theta;A_{0},B_{0},\ldots,A_{T-1},B_{T-1})$
where $I(X;Y)$ is the mutual information between random variables $X$ and $Y$.
Remark 4.4.
An important property of the Gaussian distribution is that the information gain does not depend on the observed rewards. This is because the posterior covariance of a multivariate Gaussian is a deterministic function of the sampled points. For this reason, this maximum information ratio $\gamma_{T}$ in Definition 4.3 is well defined. That is, $I(\theta;H_{T})=I(\theta;A_{0},B_{0},\ldots,A_{T-1},B_{T-1})\leqslant\gamma_{T}.$
We adopt the results from (Srinivas, Krause, Kakade, and Seeger, 2009) which gives the bounds of $\gamma_{T}$ for a range of commonly used covariance functions: finite dimensional linear, squared exponential and Matern kernels, whose details can be found in Appendix F.
Remark 4.5.
Utilizing the information of opponent’s actions and the underlying reward structure can help resolve the curse of multi-agent. For example, if the reward structure can be model by squared exponential kernels, the final regret of OTS-Hedge is $\mathcal{O}\left(\left(\sqrt{\log\mathcal{A}}+\sqrt{\log(\mathcal{A}T)\log(T)^{d+1}}\right)\sqrt{T}\right)$. In comparison to the standard adversarial multi-armed bandit regret bound $\mathcal{O}(\sqrt{T\mathcal{A}\log\mathcal{A}})$ (e.g., EXP3 do not have the information of opponent’s action), this regret bound does NOT depend polynomially on the number of actions $\mathcal{A}$ and the number of opponent’s action $\mathcal{B}$, similarly to the ideal full-information setting. This is possible due to the information of opponent action which helps capturing the underlying structure of the reward function that depends on actions of all agents.
5 Numerical Studies and Applications
Evaluation of the proposed algorithms on random matrix games and two real-world applications are conducted. IWE-Hedge (classical Exp3 algorithm) and IWE-RM serve as baseline comparisons, both adapted for bandit feedback scenarios. Moreover, Hedge and RM, which require full information, are utilized as well. Under the OTN learning framework (Algorithm 1), a total of four algorithms are compared with baselines: OTS-Hedge, OTS-RM, UCB-Hedge, and UCB-RM. Note that GP-MW (Sessa et al., 2019) is UCB-Hedge named here. Average expected regret is utilized as the performance metric. Appendix G.1 provides a detailed definition of the performance metrics and algorithm settings.
(a) Self-play
(b) Best-resp. oppo.
(c) Stat.oppo.
Figure 3: Averaged regrets for different opponents on $50\times 50$ matrix game.
5.1 Random Matrix Games
This section evaluates different algorithms on two-player zero-sum matrix games. Each payoff matrix entry is an i.i.d. random variable generated from the uniform distribution $[-1,1]$, and each player has $M$ actions (payoff matrix is a squared matrix of size $M$). The total number of rounds is $T=10^{7}$. At each round, the two players receive noisy rewards ±$\pm\tilde{r}_{t}$, where $\tilde{r}_{t}=A_{ij}+\epsilon_{t}$ and $\epsilon_{t}\sim\mathcal{N}(0,0.1)$. Different matrix sizes are considered, i.e., $M=10,50,70,100$, and for each choice of $M$, $100$ independent simulation runs are conducted. The performance, averaged over $100$ simulation runs with different opponent models (see details in Appendix G.2) are compared.
Self-play
We first compare different algorithms under the self-play setting, where both players play the same algorithm. It can be observed from Fig.2(a) that algorithms exploiting the game structure outperform the two IWE baselines, especially when the matrix size is small. Among the algorithms, OTS-based algorithms converge faster than UCB-based algorithms. Furthermore, the average regret of RM decreases earlier than that of Hedge.
Against a best-response opponent
In this setting, a best-response opponent is employed, meaning the opponent is aware of the precise value of matrix $A$ and the player’s strategy at each round. Therefore, the opponent always chooses the best action that minimizes the expected payoff. The player continues to use the various algorithms described above. It can be observed from Fig.2(b) that all the algorithms in our proposed algorithmic framework outperform the IWE baselines. Once again, the OTS-based algorithms show a faster convergence behavior compared to the UCB-based algorithms.
Against a stationary opponent
A stationary opponent is considered, whose strategy is fixed to be a probability simplex over action space. It’s important to note that the average regret in such a situation indicates the efficiency on exploiting the weakness of the opponent. Convergence curves of average regret for $M=50$ are demonstrated in Fig. 2(c). The result clearly demonstrates that the OTS estimator brings considerable benefits in exploiting this weak opponent, compared to the IWE-based estimators.
Against an non-stationary opponent
The opponent’s strategy in this subsection becomes non-stationary, and the player can only hope a robust strategy for all possible opponent’s strategies. Specifically, a game matrix $A\in\mathcal{R}^{10\times 5}$ is generated, with each element sampling from $\mathcal{N}(0.5,2.0)$. The opponent’s actions are drawn from a fixed strategy that randomly changes every $50$ rounds. Each algorithm is evaluated for up to $1000$ rounds, and $100$ simulation runs are considered. Histograms of rewards over all rounds and simulation runs are plotted in Fig. 4. The percentage of negative rewards and mean reward values are summarized in Tab. 3. Compared with IWE, OTS has smaller percentage of negative rewards and achieves higher mean rewards. This indicates OTS is more robust than IWE.
5.2 Anti-jamming Problem: Linear Game
The anti-jamming problem is an important issue in signal processing literature (Song et al., 2011), which can be formulated as a non-cooperative game between a radar and a jammer.
Figure 4: Reward histograms of different algorithms against a non-stationary opponent.
return$<0$ mean return
IWE-Hedge 19.4% 1.24
IWE-RM 12.6% 1.50
OTS-Hedge 2.5% 1.55
OTS-RM 8.8% 1.55
Table 3: Returned rewards of IWE and OTS against a non-stationary opponent.
The competition takes place at the signal level, where both parties can modify the parameters of their transmitted signals. The radar’s objective is to avoid having its transmitted signal share the same carrier frequency as the jamming signal while the jammer seeks the opposite. In this experiment, we model this competition in the frequency domain as a linear game, with the signal-to-inference-plus-noise ratio (SINR) serving as the reward function $\text{SINR}(a,b;\theta)$. Further information on the anti-jamming game setting can be found in Appendix G.7.
We compare the average regret of different algorithms against an adaptive jammer, where the jammer takes action based on the radar’s latest $10$ actions (see details in Appendix G.7). As illustrated in Fig.5, the OTS-based and UCB-based algorithms outperform the IWE-based algorithms.
(a) Hedge-based
(b) RM-based
Figure 5: Averaged regrets in the anti-jamming problem.
5.3 Repeated Traffic Routing: Kernelized Game
In this subsection, we consider the traffic routing problem from the transportation literature, which can be modeled as a multi-player game over a directed graph. Each node pair in the graph represents an individual player, and each player $i$ seeks to find the best route to send $U^{i}$ units from its origin node to its destination node. The travel time serves as the reward, depending on the traversed edges’ total occupancy. Each player’s action set is the available route in the graph, and the negative travel time of the route is the reward. A detailed description of the problem setup is provided in Appendix G.8.
We use the Sioux-Falls road network dataset (Bar-Gera, 2015) in our experiment. This network is a directed graph with $24$ nodes and $76$ edges, and there are a total of $N=528$ players. Each player $i$’s action space $\mathcal{A}^{i}$ is specified by the $5$ shortest routes, and any route that is more than three times longer than the shortest route is removed from $\mathcal{A}^{i}$. To capture the correlations in the reward function, Gaussian Process is utilized to estimate the reward function (Sessa et al., 2019). Details of the reward function estimation can be found in Appendix G.8. It is worth noting that GP-MW proposed in (Sessa et al., 2019) is equivalent to UCB-Hedge in this paper. In addition to regret, we also consider average congestion (Sessa et al., 2019) as a performance metric. Fig. 6 demonstrates that OTS-based and UCB-based algorithms outperform the IWE-based algorithms. The three proposed new algorithms, OTS-Hedge, OTS-RM, and UCB-RM, outperform UCB-Hedge (Sessa et al., 2019).
Figure 6: Average regret and congestion in traffic routing problem.
6 Conclusion
In this study, we introduced Thompson Sampling (TS)-type algorithms that utilize insights into opponents’ actions and reward structures, significantly optimizing experimental budgets with over a tenfold reduction relative to standard algorithms. Additionally, we developed the Optimism-then-NoRegret (OTN) learning framework, a versatile tool for navigating the complexities of unknown games, which encompasses a range of game algorithms as specific instances. The efficacy of our proposed solutions is validated through substantial performance enhancements observed in both synthetic scenarios and real-world applications.
Looking ahead, the convergence of emerging techniques that approximate Thompson Sampling within complex models (Li et al., 2022, 2023), such as deep neural networks, presents a fertile ground for innovation. The integration of these approaches with the OTN framework and the Optimistic Thompson Sampling (OTS) algorithm, particularly in multi-agent environments of increased complexity, holds considerable promise. Such advancements could further refine decision-making processes, enhance model robustness, and expand the applicability of our methodologies to a broader spectrum of challenges in machine learning and artificial intelligence. This prospective synergy between advanced sampling techniques and the OTN framework is poised to push the boundaries of what’s achievable in the realm of machine learning, offering exciting possibilities for future research and application.
Impact statement
Our research presents advancements in Machine Learning through the development of Thompson sampling-type algorithms for multi-agent environments with partial observations, applicable in fields like traffic routing and radar sensing. By significantly reducing experimental budgets and demonstrating a logarithmic dependence of the regret bound on the action space size, our work contributes to more efficient and equitable decision-making processes.
Ethically, this advancement promotes the responsible use of technology by ensuring broader access and reducing resource consumption. Societally, the potential applications of our research promise to enhance the functionality and safety of critical infrastructure, aligning with the goal of advancing public welfare through technological innovation. We are committed to further exploring these ethical and societal implications, ensuring that the deployment of our methodologies actively considers potential impacts to avoid unintended consequences.
In aligning with the conference’s guidelines, we recognize the established ethical impacts and societal implications common to advancements in Machine Learning. However, we have endeavored to provide a concise discussion reflective of our work’s unique contributions and potential future effects.
Appendix A Additional discussion on related works
The Nash regret defined in (O’Donoghue et al., 2021) is only meaningful when facing a best-response opponent.
Definition A.1.
Nash equilibrium and Nash value The Nash value is defind as
$V^{*}=\max_{P\in\mathcal{D}(\mathcal{A})}\min_{Q\in\mathcal{D}(\mathcal{B})}\mathbb{E}_{A\sim P,B\sim Q}\left[f_{\theta}(A,B)\right],$
and corresponding optimum $(P^{*},Q^{*})$ are the Nash equilibrium.
Definition A.2.
Nash regret in (O’Donoghue et al., 2021) The Nash regret in step $t$ is defined as
$\operatorname{NashRegret}_{t}=\mathbb{E}\left[V^{*}-R_{t+1,A_{t},B_{t}}\right]$ (7)
and the total Nash regret is defined as
$\operatorname{NashRegret}(T)=\sum_{t=1}^{T}\operatorname{NashRegret}_{t}$ (8)
We can only guarantee that $\operatorname{NashRegret}_{t}$ defined in Equation 7 is positive when the opponent is the best response player with full knowledge on the matrix, which is unrealistic. As for general opponent, $\operatorname{NashRegret}_{t}$ can be negative, which lose the meaning of ‘regret’. This is because even we can derive sublinear $\operatorname{NashRegret}(T)$, we could infer anything about the intermediate behavior of the $A$-player. A special case is that if the $A$-player can always exploit the weakness of $B$-player, the total Nash regret can be linearly decreasing to $\infty$- ∞, which is obviously ‘sublinear’. However, this exploiting situation of the two players should be distinctive to the case that two players are playing Nash equilibrium and resulting 00 total Nash regret, which is also ‘sublinear’.
As for our definition of adversarial regret in Equation 1 in this repeated game setting with unknown reward function, we expect to measure how the $A$-player could exploit the weakness of $B$-player as time going on. The negative regret would not appear in any cases of the opponent. This is our motivation to use adversarial regret.
Appendix B Description of Games and Full Information Game
In this work, we consider three representative game forms: the matrix game, linear game, and kernelized game, as summarized in Table 4.
Table 4: Game Examples.
Matrix Game Linear Game Kernelized Game
Mean Reward $f_{\theta}(a,b)=\theta_{a,b}$ $f_{\theta}(a,b)=\phi(a,b)^{\top}\theta$ $f_{\theta}(a,b)=\theta(a,b)$
$\theta\in\mathbb{R}^{\mathcal{A}\times\mathcal{B}}$ $\phi(a,b)\in\mathbb{R}^{d}$, $\theta\in\mathbb{R}^{d}$ $\theta(a,b)\in\mathbb{R}$ is mean of a Gaussian Process
B.1 Different Games
Assumption B.1.
The corruption noise $W_{t}=Y_{t+1,A_{t},B_{t}}-f_{\theta}(A_{t},B_{t})$ is assumed to be zero-mean Gaussian noise and is independent at each time $t$.
Example B.1 (Matrix games).
In a matrix game, the reward function simplifies to $f_{\theta}(a,b)=\theta_{a,b}$. In this degenerate setting, $\theta$ can be considered as the utility matrix for Alice.
Example B.2 (Linear games).
In a linear game, a known feature mapping $\phi:\mathcal{A}\times\mathcal{B}\mapsto\mathbb{R}^{d}$is defined, and the mean reward function is given by $f_{\theta}(a,b)=\phi(a,b)^{\top}\theta$, where the reward is linear in the feature. We assume that the random parameter $\theta$ follows the normal distribution ${N}(\mu_{p},\Sigma_{p})$, and the reward noise $W_{t+1}=Y_{t+1,A_{t},B_{t}}-f_{\theta}(A_{t},B_{t})$ is normally distributed with with mean zero and variance $\sigma_{w}^{2}$, independent of $(H_{t},A_{t},B_{t},\theta)$.
In the Section 5.3, the reward structure in the repeated traffic routing problem is modeled using a kernel function, which is referred to as Kernelized games.
Example B.3 (Kernelized games).
In kernelized games, we consider the case where the reward function $f_{\theta}$ is sample from a Gaussian process. The stochastic process $(f_{\theta}(a,b):(a,b)\in\mathcal{A}\times\mathcal{B})$ follows a multivariate Gaussian distribution, where the mean function is denoted as $\mu(a,b)=\mathbb{E}\left[f_{\theta}(a,b)\right]$ and covariance (or kernel) function is denoted as $k((a,b),(a^{\prime},b^{\prime}))=\mathbb{E}\left[(f_{\theta}(a,b)-\mu(a,b))(f_{\theta}(a^{\prime},b^{\prime})-\mu(a^{\prime},b^{\prime}))\right]$. The kernel function $k((a,b),(a^{\prime},b^{\prime}))$ measures the similarity between different action pairs $(a,b),(a^{\prime},b^{\prime})\in\mathcal{A}\times\mathcal{B}$ in the game. We assume that the function $f_{\theta}$ is sampled from a Gaussian process prior $GP(0,k((a,b),(a^{\prime},b^{\prime})))$, the reward noise $W_{t+1}=Y_{t+1,A_{t},B_{t}}-f_{\theta}(A_{t},B_{t})$ is independent of $(H_{t},\theta,A_{t},B_{t})$, and $(W_{t}:t\in\mathbb{Z}_{++})$ is an i.i.d sequence following $N(0,\sigma_{w}^{2})$.
B.2 Full Information Feedback
To introduce the proposed Optimism-then-NoRegret learning framework, we first consider the full information feedback setting where Alice can observe the mean rewards $r_{t}(a)=f_{\theta}(a,B_{t})$ for all actions $a\in\mathcal{A}$. In this case, the problem can be solved using full-information adversarial bandit algorithms such as Hedge (Freund & Schapire, 1997) and Regret Matching (RM) (Hart & Mas-Colell, 2000) applied to the sequence of adversarial reward vectors $(r_{t})_{t\in[T]}\in[0,1]^{\mathcal{A}\times T}$. The procedure is summarized in Algorithm 2, where $P_{X}$ denotes probability simplex proportional to $X$, and function $g_{t}:\Delta^{\mathcal{A}}\times[0,1]^{\mathcal{A}}\mapsto\Delta^{\mathcal{A}}$ in round $t$ is specified as follows:
[…]
In the full information setting, where $r_{t}(a)=f_{\theta}(a,B_{t})$, the adversarial regret defined in Equation 1 translates to the following full information adversarial regret:
Definition B.2 (Regret with Full Information).
The full information adversarial regret of algorithm $\operatorname{adv}$ for arbitrary reward sequence $(r_{t})_{t}$ is defined as
$\Re_{\operatorname{full}}(T,{\operatorname{adv}},(r_{t})_{t})=\max_{a\in\mathcal{A}}\mathbb{E}\left[\sum_{t=0}^{T-1}r_{t}(a)-r_{t}(A_{t})\right]$ (9)
The following proposition provides the regret bounds for Hedge and RM algorithms in the full information feedback setting (Freund & Schapire, 1997):
Proposition B.3 (Regrets of Hedge and RM).
Consider playing in a full information feedback game with Hedge or RM algorithms. The regrets are bounded as follows:
$\Re_{\operatorname{full}}(T,{\operatorname{Hedge}},(r_{t})_{t})=\mathcal{O}(\sqrt{T\log\mathcal{A}}),\ \Re_{\operatorname{full}}(T,{\operatorname{RM}},(r_{t})_{t})=\mathcal{O}(\sqrt{T\mathcal{A}}).$
Algorithm 2 No Regret Update for Full-Information Feedback
1:  Initialize $X_{1}$
2:  for round $t=0,1,\ldots,T-1$  do
3:     Sample action $A_{t}\sim P_{X_{t}}$,
4:     Observe full-information feedback $f_{\theta}(a,B_{t})$ for all $a\in\mathcal{A}$,
5:     Update: $X_{t+1}=g_{t}(X_{t},(f_{\theta}(a,B_{t}))_{a\in\mathcal{A}})$.
6:  end for
Appendix C Updating Rules of Different Games
The B.1 together with the stucture of mean reward functions gives the following Bayesian update rule of posterior. It is also possible to use Bayesian updated algorithm in frequentist setting, with a slightly different treatment.
Posterior distribution for Linear Gaussian model (Parametric).
Let’s consider the linear Gaussian model with a Gaussian prior $N(\mu_{p},\Sigma_{p})$ and noise likelihood $N(0,\sigma_{w}^{2})$. Here are the key aspects of the model:
  • Prior Assumptions: We assume a zero-mean prior with covariance $\Sigma_{p}=\sigma_{p}I$, where $\sigma_{p}$ is a scalar parameter satisfying $\sigma_{p}\leqslant 1$.
  • Feature Map Assumptions: We assume that the feature map $\phi(a,b)$ satisfies $|\phi(a,b)|\leqslant 1$.
  • Covariance Matrix Update: Given the initial covariance matrix $\Sigma_{0}=\Sigma_{p}$, the covariance matrix at time $t+1$ is updated as:
    $\Sigma_{t+1}=\left(\Sigma_{t}^{-1}+\frac{1}{\sigma_{w}^{2}}\phi(A_{t},B_{t})\phi(A_{t},B_{t})^{\top}\right)^{-1}$
  • Mean Vector Update: Given the initial mean vector $\mu_{0}=\mu_{p}$, the mean vector at time $t+1$ is updated as:
    $\mu_{t+1}=\Sigma_{t+1}\left(\Sigma_{t}^{-1}\mu_{t}+\frac{R_{t+1,A_{t},B_{t}}}{\sigma_{w}^{2}}\phi(A_{t},B_{t})\right)$
  • $\sigma_{t}(a,b)=\left\|\phi(a,b)\right\|_{\Sigma_{t}}$ and $\mu_{t}(a,b)=\phi(a,b)^{\top}\mu_{t}$
  • In the case where the feature $\phi(a,b)=e_{a,b}$ is a one-hot vector, we denote $n_{t}(a,b)$ as the counts of occurrences of $(a,b)$ up to time $t$, the posterior variance is given by:
    $\sigma_{t}(a,b)=\sqrt{\frac{\sigma_{w}^{2}}{\sigma_{w}^{2}/\sigma^{2}_{p}(a,b)+n_{t}(a,b)}}$
Posterior distribution for Gaussian Process (Non-paramatric).
In the non-parametric case of a Gaussian Process (GP), the posterior distribution remains Gaussian as well. Here are the relevant details:
  • Notation: We define the vector $\mathbf{k}_{t}((a,b))$ and $\mathbf{R}_{t}$, and the matrix $\mathbf{K}_{t}$ as follows:
    $\mathbf{k}_{t}(a,b)$ $=[k((A_{0},B_{0}),(a,b)),\ldots,k((A_{t-1},B_{t-1}),(a,b))]^{\top}$
    $\mathbf{R}_{t}$ $=[R_{1,A_{0},B_{0}},\ldots,R_{t,A_{t-1},B_{t-1}}]^{\top}$
    $\mathbf{K}_{t}(i,j)$ $=k((A_{i},B_{i}),(A_{j},B_{j}))$
  • Variance Assumption: We assume that the variance satisfies $k(x,x)\leqslant 1$ for all $x\in\mathcal{X}$.
  • Posterior Variance: The posterior variance at time $t$ is given by: $\sigma^{2}_{t}(a,b)=k((a,b),(a,b))-\mathbf{k}_{t}((a,b))^{\top}(\mathbf{K}_{t}+\sigma^{2}{\bm{I}}_{t})\mathbf{k}_{t}(a,b)$
  • Posterior Mean: The posterior mean at time $t$ is given by: $\mu_{t}(a,b)=\mathbf{k}_{t}((a,b))^{\top}(\mathbf{K}_{t}+\sigma^{2}{\bm{I}}_{t})^{-1}\mathbf{R}_{t}$
  • Relationship to Linear Gaussian Model: If the kernel $k((a,b),(a,b))=\phi(a,b)^{\top}\Sigma_{p}\phi(a,b)$ is composed of basis functions, the GP reduces to the linear Gaussian model with a prior covariance matrix $\Sigma_{p}$. This ensures coherence between the linear Gaussian model and the kernel model assumptions.
Appendix D Details for Importance Weighted Estimator in Section 3
Importance-weighted estimator.
For any measurable function $h$ and probability distribution $(X_{a})_{a\in\mathcal{A}}$ over a finite support $\mathcal{A}$, we construct importance weighted estimator
$\tilde{h}(a)=\mathbb{I}_{A=a}\frac{h(A)}{X_{a}},\forall a\in\mathcal{A}$
which is an unbiased estimator:
$\mathbb{E}\left[\tilde{h}(a)\right]=\mathbb{E}\left[\mathbb{I}_{A=a}\right]h(a)/X_{a}=h(a).$
Exp3: Hedge with importance weighted estimator
In this work, we sometimes call Exp3 as IWE-Hedge. The celebrated Exp3 algorithm construct an estimate of reward vector as
$\tilde{R}_{t+1}(a)=1-\frac{\mathbb{I}_{A_{t}=a}(1-R_{t+1,A_{t},B_{t}})}{X_{t,a}}$
We can observe that $\tilde{R}_{t+1}(a)$ is unbiased conditioned on history $H_{t}$. Given that $X_{t}$ is $H_{t}$-measurable and $A_{t}$ is conditionally independent with $W_{t+1}$ and $B_{t}$ given $H_{t}$, and using the fact $\mathbb{I}_{A_{t}=a}R_{t+1,A_{t},B_{t}}=\mathbb{I}_{A_{t}=a}R_{t+1,a,B_{t}}$, we have:
$\mathbb{E}_{t}\left[\tilde{R}_{t+1}(a)\right]=1-\mathbb{E}_{t}\left[\mathbb{I}_{A_{t}=a}\frac{1-R_{t+1,a,B_{t}}}{X_{t,a}}\right]=1-\mathbb{E}_{t}\left[\mathbb{I}_{A_{t}=a}\right]\frac{1-\mathbb{E}_{t}\left[f_{\theta}(a,B_{t})\right]}{X_{t,a}}=\mathbb{E}_{t}\left[f_{\theta}(a,B_{t})\right].$
Exp3(Auer et al., 2002b) updates the strategy using $X_{t+1,a}\propto X_{t,a}\exp(\eta_{t}\tilde{R}_{t+1}(a))$.
Regret Matching with importance weighted estimator.
Using the importance-weighted estimator, we can obtain an unbiased estimator for the regret $\Re_{t}\in\mathbb{R}^{\mathcal{A}}$ at round $t$:
~$\tilde{\Re}_{t,a}=\frac{\mathbb{I}_{A_{t}=a}R_{t+1,A_{t},B_{t}}}{X_{t,a}}-R_{t+1,A_{t},B_{t}}\frac{\hat{X}_{t,A_{t}}}{X_{t,A_{t}}}$ (10)
and update the strategy as follows: Let the cumulative estimated reward be $\tilde{C}_{t,a}=\sum_{s=0}^{t}\tilde{\Re}_{s,a}$,
$\hat{X}_{t+1,a}=\begin{cases}{\tilde{C}^{+}_{t,a}}/{\sum_{a\in\mathcal{A}}\tilde{C}_{t,a}^{+}},&\text{if }\sum_{a\in\mathcal{A}}\tilde{C}_{t,a}^{+}>0,\\ \text{arbitrary vector on simplex, e.g. }1/\mathcal{A},&\text{otherwise}\end{cases}$ (11)
Here, the sampling distribution $X_{t,a}$ is mixed with uniform distribution
$X_{t,a}=(1-\gamma_{t})\hat{X}_{t,a}+\gamma_{t}(1/\mathcal{A}),\forall a\in\mathcal{A}$ (12)
The detailed algorithm for Importance-weighted estimator Regret Matching (IWE-RM) is as follows:
Algorithm 3 Importance-weighted estimator with Regret Matching (IWE-RM)
1:  Input: init $X_{1}=\hat{X}_{1}$ as uniform probability vector over $\mathcal{A}$ and sequence $(\gamma_{t})_{t\geqslant 0}$
2:  for round $t=0,1,\ldots,T-1$  do
3:     Sample action $A_{t}\sim P_{X_{t}}$
4:     Observe noisy bandit feedback $R_{t+1,A_{t},B_{t}}$.
5:     Construct regret estimator ~$\tilde{\Re}_{t}$ with importance weighted estimation by Equation 10.
6:     Update $\hat{X}_{t+1}$ and $X_{t+1}$ with Equations 11 and 12 and $\gamma_{t}$.
7:  end for
Fact D.1.
Importance weighted estimator $\tilde{R}_{t+1}$ at round $t$ is $\sigma(H_{t},A_{t},R_{t+1,A_{t},B_{t}})$-measurable.
Remark D.1.
For any $t\in\mathbb{N}$, the imagined reward vector $\tilde{R}_{t+1}$ constructed by importance weighted estimator satisfies $\mathbb{E}\left[\tilde{R}_{t+1}(a)\>|\>H_{t},\theta\right]=\mathbb{E}\left[f_{\theta}(a,B_{t})\>|\>H_{t},\theta\right]=\mathbb{E}\left[g_{\theta}(e_{a},Y_{t})\>|\>H_{t},\theta\right]$ for any $a\in\mathcal{A}$. Therefore, we have $\mathbb{E}\left[\operatorname{pess}_{t+1}\>|\>\theta\right]=\mathbb{E}\left[\operatorname{est}_{t+1}\>|\>\theta\right]=0$.
Lemma D.2.
For all real $a$, define $a^{+}=\max\{a,0\}$. For all $a,b$, it is the case that
$\left((a+b)^{+}\right)^{2}\leqslant(a^{+})^{2}+2(a^{+})b+b^{2}$
Proof.
$(a+b)^{+}\leqslant(a^{+}+b)^{+}\leqslant\left|a^{+}+b\right|.$
Lemma D.3.
For all vector $v\in\mathbb{R}^{\mathcal{A}}$, define $v^{+}=(v_{a}^{+})_{a\in\mathcal{A}}$. Following Algorithm 3, we have the important observation
$\left\langle\tilde{C}_{t-1}^{+},\tilde{\Re}_{t}\right\rangle\leqslant 0$
Proof.
Suppose at round $t$, Alice choose $A_{t}\sim P_{X_{t}}$ and receive the feedback $R_{t+1,A_{t},B_{t}}$. By algorithm 3 and Equations 10, 11 and 12, If $\sum_{a}\tilde{C}_{t-1,a}^{+}\leqslant 0$, then obviously $\tilde{C}^{+}_{t-1,a}=0$ for all action $a\in\mathcal{A}$. Then, the lemma trivially holds.
Otherwise, we have
$\left\langle\tilde{C}_{t-1}^{+},\tilde{\Re}_{t}\right\rangle$ $=R_{t+1,A_{t},B_{t}}\left(\frac{\tilde{C}_{t-1,A_{t}}^{+}}{X_{t,A_{t}}}-\frac{\hat{X}_{t,A_{t}}}{{X}_{t,A_{t}}}\sum_{a}\tilde{C}_{t-1,a}^{+}\right)$
$=R_{t+1,A_{t},B_{t}}\left(\frac{\tilde{C}_{t-1,A_{t}}^{+}}{X_{t,A_{t}}}-\frac{\tilde{C}_{t-1,A_{t}}^{+}/\sum_{a}\tilde{C}_{t-1,a}^{+}}{{X}_{t,A_{t}}}\sum_{a}\tilde{C}_{t-1,a}^{+}\right)=0$
Lemma D.4.
Following algorithm 3, we have an important inequality
$\sum_{a}\left({\tilde{C}_{T,a}^{+}}\right)^{2}\leqslant\sum_{t=1}^{T}\sum_{a}\left(\tilde{\Re}_{t,a}\right)^{2}$
Proof.
Since from the update rule,
${\tilde{C}_{T,a}}={\tilde{C}_{T-1,a}}+\tilde{\Re}_{T,a},$
by Lemma D.2 and D.3:
$\sum_{a}\left({\tilde{C}_{T,a}^{+}}\right)^{2}$ $\leqslant\sum_{a}\left(\left({\tilde{C}_{T-1,a}^{+}}\right)^{2}+2{\tilde{C}_{T-1,a}^{+}}{\tilde{\Re}_{T,a}}+\left(\tilde{\Re}_{T,a}\right)^{2}\right)$
$\leqslant\sum_{a}\left({\tilde{C}_{T-1,a}^{+}}\right)^{2}+\sum_{i}(\tilde{\Re}_{T,a})^{2}$
By telescoping series, we conclude the lemma. ∎
Proof of Theorem 3.3.
Notice that the action $A_{t}$ selected by Alice and the action $B_{t}$ selected by Bob is independent conditioned on history $H_{t}$ and $\theta$. According to the algorithm 3, the conditional expectation of the estimated immediate regret is
$\mathbb{E}\left[\tilde{\Re}_{t,a}\mid\theta,H_{t}\right]=\mathbb{E}\left[f_{\theta}(a,B_{t})\mid\theta,H_{t}\right]-\sum_{a}\hat{X}_{t,a}\mathbb{E}\left[f_{\theta}(a,B_{t})\mid\theta,H_{t}\right]$
Step 1 (Bounding the bias of estimated regret.)
Recall the definition of immediate regret at time $t$ conditioned on history is
$\Re_{t}(a):=\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\mid\theta,H_{t}\right]=\mathbb{E}\left[f_{\theta}(a,B_{t})\mid\theta,H_{t}\right]-\sum_{a}{X}_{t,a}\mathbb{E}\left[f_{\theta}(a,B_{t})\mid\theta,H_{t}\right]$
In the following, we use short notation $\mathbb{E}_{t}\left[\cdot\right]=\mathbb{E}\left[\cdot\mid H_{t},\theta\right]$. For any $a\in\mathcal{A}$, the difference with the estimated regret under conditional expectation is
$\Re_{t}(a)-\mathbb{E}_{t}\left[\tilde{\Re}_{t,a}\right]$ $=\sum_{a}(\hat{X}_{t,a}-X_{t,a})\sum_{b}Y_{t,b}f_{\theta}(a,b)$
$=\sum_{a}(\gamma_{t}\hat{X}_{t,a}-\gamma_{t}/\mathcal{A})\sum_{b}Y_{t,b}f_{\theta}(a,b)$
$\leqslant\sum_{a}\left|\gamma_{t}\hat{X}_{t,a}-\gamma_{t}/\mathcal{A}\right|$
$\leqslant\sum_{a}\left(\left|\gamma_{t}\hat{X}_{t,a}\right|+\left|\gamma_{t}/\mathcal{A}\right|\right)=2\gamma_{t}$
Step 2 (Bounding the potential.)
For any $a\in\mathcal{A}$,
$\mathbb{E}_{0}\left[\tilde{C}_{T,a}\right]$ $\leqslant\mathbb{E}_{0}\left[\tilde{C}_{T,a}^{+}\right]=\mathbb{E}_{0}\left[\sqrt{(\tilde{C}_{T,a}^{+})^{2}}\right]\leqslant\mathbb{E}_{0}\left[\sqrt{\sum_{a}\left(\tilde{C}_{T,a}^{+}\right)^{2}}\right],$
where the last inequality is due to the Jensen inequality. By lemma D.4 and taking expectation,
$\mathbb{E}_{0}\left[\sqrt{\sum_{a}\left(\tilde{C}^{+}_{T,i}\right)^{2}}\right]\leqslant\mathbb{E}_{0}\left[\sqrt{\sum_{t=1}^{T}\sum_{a}\left(\tilde{\Re}_{t,a}\right)^{2}}\right]$
The RHS of the above inequality can be bounded as
$\mathbb{E}_{0}\left[\sum_{t=1}^{T}\sum_{a}\left(\tilde{\Re}_{t,a}\right)^{2}\right]$ $=\mathbb{E}_{0}\left[\sum_{t=1}^{T}\sum_{a}\left(\frac{\mathbb{I}_{A_{t}=a}R_{t+1,A_{t},B_{t}}}{X_{t,a}}-R_{t+1,A_{t},B_{t}}\frac{\hat{X}_{t,A_{t}}}{X_{t,A_{t}}}\right)^{2}\right]$
$=\mathbb{E}_{0}\left[\sum_{t=1}^{T}\sum_{a}R_{t+1,A_{t},B_{t}}^{2}\left(\left(\frac{\mathbb{I}_{A_{t}=a}}{X_{t,A_{t}}}\right)^{2}-\frac{2\hat{X}_{t,A_{t}}}{X_{t,A_{t}}^{2}}\mathbb{I}_{A_{t}=a}+\frac{\hat{X}_{t,A_{t}}^{2}}{X_{t,A_{t}}^{2}}\right)\right]$
$=\mathbb{E}_{0}\left[\sum_{t=1}^{T}R_{t+1,A_{t},B_{t}}^{2}\left(\frac{1}{X_{t,A_{t}}^{2}}-\frac{2\hat{X}_{t,A_{t}}}{X_{t,A_{t}}^{2}}+\left|\mathcal{A}\right|\frac{\hat{X}_{t,A_{t}}^{2}}{X_{t,A_{t}}^{2}}\right)\right]$
$=\mathbb{E}_{0}\left[\sum_{t=1}^{T}\mathbb{E}_{t}\left[R_{t+1,A_{t},B_{t}}^{2}\left(\frac{1}{X_{t,A_{t}}^{2}}-\frac{2\hat{X}_{t,A_{t}}}{X_{t,A_{t}}^{2}}+\left|\mathcal{A}\right|\frac{\hat{X}_{t,A_{t}}^{2}}{X_{t,A_{t}}^{2}}\right)\right]\right],$
where we have the following derivation by the fact
$\mathbb{E}_{t}\left[R_{t,a,b}^{2}\right]:=\mathbb{E}_{t}\left[(f_{\theta]}(a,b)+W_{t+1})^{2}\right]=\mathbb{E}_{t}\left[f_{\theta]}(a,b)^{2}+W_{t+1}^{2}\right]\leqslant 1+\sigma_{w}^{2}$
and the fact $0\leqslant{\hat{X}_{t,a}}/{X}_{t,a}\leqslant 1/(1-\gamma_{t})$ for all $a\in\mathcal{A}$,
$\mathbb{E}_{t}\left[R_{t+1,A_{t},B_{t}}^{2}\left(\frac{1}{X_{t,A_{t}}^{2}}-\frac{2\hat{X}_{t,A_{t}}}{X_{t,A_{t}}^{2}}+\left|\mathcal{A}\right|\frac{\hat{X}_{t,A_{t}}^{2}}{X_{t,A_{t}}^{2}}\right)\right]$
$=\sum_{a}\frac{\mathbb{E}_{t}\left[R_{t+1,a,B_{t}}^{2}\right]}{X_{t,a}}+\sum_{a}\frac{\hat{X}_{t,a}}{{X}_{t,a}}\mathbb{E}_{t}\left[R_{t+1,A_{t},B_{t}}^{2}\right]\left(\left|\mathcal{A}\right|\hat{X}_{t,a}-2\right)$
$\leqslant(1+\sigma_{w}^{2})\left(\sum_{a}\frac{1}{{X}_{t,a}}+\sum_{a}\frac{\hat{X}_{t,a}}{{X}_{t,a}}\left(\left|\mathcal{A}\right|\hat{X}_{t,a}-2\right)\right)$
$\leqslant(1+\sigma_{w}^{2})\left(\sum_{a}\frac{1}{X_{t,a}}+\sum_{a}\frac{\hat{X}_{t,a}}{{X}_{t,a}}\left(\left|\mathcal{A}\right|-2\right)\right)$
$\leqslant(1+\sigma_{w}^{2})\left(\sum_{a}\frac{\left|\mathcal{A}\right|}{\gamma_{t}}+\min(\frac{\left|\mathcal{A}\right|}{\gamma_{t}},\frac{\left|\mathcal{A}\right|}{1-\gamma_{t}})(\left|\mathcal{A}\right|-2)\right)\leqslant\frac{2(1+\sigma_{w}^{2})\left|\mathcal{A}\right|^{2}}{\gamma_{t}}$
Then, we derive one important relationship
$\mathbb{E}_{0}\left[\sqrt{\sum_{a}\left(\tilde{C}^{+}_{T,i}\right)^{2}}\right]\leqslant\mathbb{E}_{0}\left[\sqrt{\sum_{t=1}^{T}\sum_{a}\left(\tilde{\Re}_{t,a}\right)^{2}}\right]\leqslant\sqrt{\sum_{t=1}^{T}\frac{2(1+\sigma_{w}^{2})\left|\mathcal{A}\right|^{2}}{\gamma_{t}}}$
Step 3 (Put all together.)
$\mathbb{E}_{0}\left[\sum_{t=1}^{T}\Re_{t}(a)\right]=\mathbb{E}_{0}\left[\sum_{t=1}^{T}\mathbb{E}_{t}\left[\tilde{\Re}_{t,a}\right]+\sum_{t=1}^{T}2\gamma_{t}\right]=\mathbb{E}_{0}\left[\tilde{C}_{T,a}+\sum_{t=1}^{T}2\gamma_{t}\right]\leqslant\sqrt{\sum_{t=1}^{T}\frac{2(1+\sigma_{w}^{2})\left|\mathcal{A}\right|^{2}}{\gamma_{t}}}+2\gamma_{t}$
When $\gamma_{t}=\gamma$,
$\mathbb{E}_{0}\left[\sum_{t=1}^{T}\Re_{t}(a)\right]\leqslant\sqrt{T}\sqrt{\frac{2(1+\sigma_{w}^{2})\left|\mathcal{A}\right|^{2}}{\gamma}}+2\gamma T,$
Taking $\gamma=\sqrt[3]{((1+\sigma_{w}^{2})\left|\mathcal{A}\right|^{2})/{2T}}$, we have
$\mathbb{E}_{0}\left[\sum_{t=1}^{T}\Re_{t}(a)\right]\leqslant 2^{4/3}(1+\sigma_{w}^{2})^{1/3}\left|\mathcal{A}\right|^{2/3}T^{2/3}.$
Appendix E Analysis of the failure for Thompson sampling estimator in Section 3.3
E.1 Basic setting of the counter example
Consider a class of matrix games with a payoff matrix $\theta$ defined as
$\theta=\begin{bmatrix}1&1-\Delta\\ 1-\Delta&1\end{bmatrix},$ (13)
where $\Delta\in(0,1)$. At time $t$, Alice plays action $A_{t}\sim P_{X_{t}}$, and Bob is a best response player who can observe $X_{t}$ and play the best response strategy $Y_{t}=\max_{y}X_{t}^{\top}\theta y$ by selecting action $B_{t}\sim P_{Y_{t}}$. Alice receives the noiseless reward $R_{t+1,A_{t},B_{t}}=\theta_{A_{t},B_{t}}$. From the example $\theta$, we can observe the following:
  • Observation $1$: When Alice uses a pure strategy, she suffers a regret of $\Delta$ at that round due to Bob’s best-response strategy.
  • Observation $2$: The best-response strategy for a uniform strategy is also a uniform strategy.
Let’s define the following terms:
  • $X_{t}$ and $reg_{t}$: Alice’s strategy and instantaneous regret at time $t$.
  • $\tilde{R}_{t}$ and $m_{t}$: $\tilde{R}_{t}=\left[\tilde{R}_{t1},\tilde{R}_{t2}\right]$ is the estimated reward vector by TS-RM estimator; $m_{t}=\tilde{R}_{t1}-\tilde{R}_{t2}$ represents the difference between two rewards.
Now, let’s consider the following remark regarding initialization:
Remark E.1 (Initialization).
The TS-RM algorithm for Alice, initialized with a uniform strategy, will always result in a pure strategy in $X_{2}$.
Proof.
In the regret-matching algorithm, the instantaneous regret $reg_{1}$ at $t=1$ can be represented as
$reg_{1}=\tilde{R}_{1}-\tilde{R}_{1}^{T}X_{1}\cdot\mathbf{1},$ (14)
Since Alice and Bob are both initialized with a uniform strategy, i.e., $X_{1}=[0.5,0.5]$, it can be observed that if $m_{1}\neq 0$, the two elements in $reg_{1}$ will always have opposite signs. If $m_{1}=0$, $X_{2}=X_{1}$ and can still be regarded as an initialization step.According to the regret-matching updating rule, $X_{2}\propto\max(reg_{1},0)$, which means $X_{2}$ must be a pure strategy. ∎
Based on Remark E.1, we can draw the following conclusions regarding the counter example:
  • The choice of uniform initialization for the TS-RM algorithm does not affect the divergence result.
  • This result holds regardless of the specific action chosen by Alice at $t=1$, indicating that two symmetric conditions arise depending on whether $X_{2}=[1,0]$ or $X_{2}=[0,1]$.
E.2 TS-RM suffers linear regret
According to Remark E.1, without loss of generality (w.l.o.g.), let us define the following events:
  • Event $\omega_{t}$: Alice picks the second row and the best-response opponent chooses the first column at time $t$.
  • Event $\Omega_{t}$: Alice picks the second row and the best-response opponent chooses the first column for all time $t^{\prime}\leqslant t$.
The occurrence of event $\Omega_{t}$ implies that Alice experiences linear regret until time $t$. However, the actual convergence probability is greater than $\mathbb{P}(\Omega_{t})$ since even if $\Omega_{t}$ does not occur (i.e., Alice occasionally chooses the optimal result $1$), there is still a probability that TS-RM fails. Quantifying this probability is challenging. If we can demonstrate that $\Omega_{t}$ occurs with a constant probability $c$, then the divergence probability of TS-RM should be greater than $c$. Specifically, due to the symmetric property of the example $\theta$, we obtain the following propositions:
Proposition E.2.
If Alice initializes with a uniform strategy,
$\Re(T)\geqslant 2\mathbb{P}(\Omega_{T})\Delta\cdot T$
If $\Omega_{t}$ happens with constant probability for all $t\geqslant 1$, then Alice suffers linear regret.
Proposition E.3 (Failure of TS-RM).
Suppose Alice initializes with a uniform strategy and utilizes Regret Matching with Thompson Sampling estimator (TS-RM). For any $\Delta\in(0,1)$ and $\sigma_{n}>0$, there exists a constant $c(\Delta,\sigma_{n})>0$ such that for all rounds $t\geqslant 1$, $\mathbb{P}(\Omega_{t})\geqslant c(\Delta,\sigma_{n})$.
To investigate the divergence behavior of the TS-RM algorithm, an experiment is conducted using the counter example in . Different values of $\Delta$ and $\sigma_{n}^{2}$ are considered, and $200$ independent simulation runs are performed for each combination. The averaged divergence results across these runs are shown in Fig.7, which illustrates that the probability of divergence decreases as $\Delta$ and $\sigma_{n}^{2}$ increase, consistent with our proposition (Prop.E.3). In the following, we provide a detailed proof for the divergence of the TS-RM algorithm.
Figure 7: Divergence probability for TS-RM
Proof.
Our goal is to prove that there exists a constant $c\geqslant 0$ such that $\mathbb{P}(\Omega_{t})\geqslant c$ for all $t$. The probability $\mathbb{P}(\Omega_{t})$ can be expressed as
$\mathbb{P}(\Omega_{t})=\mathbb{P}(\omega_{1})\mathbb{P}(\omega_{2}|\Omega_{1})\ldots\mathbb{P}(\omega_{t}|\Omega_{t-1})$ (15)
Referring to Remark E.1, $X_{2}$ is a pure strategy. Since Alice chooses the second row according to $\Omega_{2}$, we get $X_{2}=[0,1]$ and $m_{1}<0$. Following the event $\Omega_{t}$, we get $X_{t}=[0,1],\forall t\geqslant 2$, which indicates that for the TS-RM estimator $\tilde{R}_{t}$, only the posterior distribution of $\tilde{R}_{t2}$ is updated. Since no noise is considered in received rewards, by Bayesian rule, we have
$\tilde{R}_{t}=\left[z_{t},\frac{t}{t+\sigma_{n}^{2}}(1-\Delta)+\sqrt{\frac{\sigma_{n}^{2}}{\sigma_{n}^{2}+t}}z^{\prime}_{t}\right],$
where $z_{t},z^{\prime}_{t}\sim\mathcal{N}(0,1)$ are two independent r.v.s, and $\sigma_{n}^{2}$ is the noise variance. Therefore, we can express the regret as:
$reg_{t}=\tilde{R}_{t}-X_{t}^{T}\tilde{R}_{t}\cdot\mathbf{1}=[m_{t},0],\quad\forall t\geqslant 2$ (16)
where $m_{t}=z_{t}-\frac{t}{t+\sigma_{n}^{2}}(1-\Delta)-\sqrt{\frac{\sigma_{n}^{2}}{t+\sigma_{n}^{2}}}z^{\prime}_{t}=z_{t}-\sqrt{\frac{\sigma_{n}^{2}}{t+\sigma_{n}^{2}}}z^{\prime}_{t}-\frac{t}{t+\sigma_{n}^{2}}(1-\Delta)$.
The cumulative regret can be represented as
$\Re_{t}=[0.5m_{1}+\sum_{k=2}^{t}m_{k},-0.5m_{1}]$ (17)
According to updating rule of TS-RM, we have:
$\mathbb{P}(\omega_{t}|\Omega_{t-1})$ $=\mathbb{P}(0.5m_{1}+\sum_{k=2}^{t}m_{k}\leqslant 0|\Omega_{t-1})$ (18)
$\geqslant\mathbb{P}(\sum_{k=2}^{t}m_{k}\leqslant 0|\Omega_{t-1})$
$=\mathbb{P}(\sum_{k=2}^{t}m_{k}\leqslant 0|\Omega_{t-1})\left(\mathbb{P}(\Omega_{t-1})+\mathbb{P}(\bar{\Omega}_{t-1})\right)$
$\geqslant\mathbb{P}(\sum_{k=2}^{t}m_{k}\leqslant 0|\Omega_{t-1})\mathbb{P}(\Omega_{t-1})+\mathbb{P}(\sum_{k=2}^{t}m_{k}\leqslant 0|\bar{\Omega}_{t-1})\mathbb{P}(\bar{\Omega}_{t-1})$
$=\mathbb{P}(\sum_{k=2}^{t}m_{k}\leqslant 0)$
where the first inequality is because $m_{1}\leqslant 0$ (conditioned on $\Omega_{t-1}$), and the second one is due to $\mathbb{P}(\sum_{k=2}^{t}m_{k}\leqslant 0|\Omega_{t-1})\geqslant\mathbb{P}(\sum_{k=2}^{t}m_{k}\leqslant 0|\bar{\Omega}_{t-1})$.
Define
$N_{t}=\sum\limits_{k=2}^{t}m_{k}\sim\mathcal{N}\left(-\sum\limits_{k=2}^{t}\frac{k}{k+\sigma_{n}^{2}}(1-\Delta),\ \sum\limits_{k=2}^{t}(1+\frac{\sigma_{n}^{2}}{k+\sigma_{n}^{2}})\right)\triangleq\mathcal{N}(\mu_{t},\sigma_{t}^{2}).$
As $t\to\infty$, we have:
$\lim_{t\to\infty}\log\mathbb{P}(\Omega_{t})$ $=\log\prod_{t=1}^{\infty}\mathbb{P}(\omega_{t}|\Omega_{t-1})=\sum_{t=1}^{\infty}\log\mathbb{P}(\omega_{t}|\Omega_{t-1})$ (19)
$\geqslant\sum_{t=1}^{\infty}\log\mathbb{P}(N_{t}\leqslant 0)$
$=\sum_{t=1}^{\infty}\log\mathbb{P}(\mu_{t}+\sigma_{t}Z\leqslant 0),\quad Z\sim\mathcal{N}(0,1)$
$=\sum_{t=1}^{\infty}\log\mathbb{P}(Z\leqslant-\frac{\mu_{t}}{\sigma_{t}})$
$=\sum_{t=1}^{\infty}\log\Phi(-\frac{\mu_{t}}{\sigma_{t}})$
$=\sum_{t=1}^{\infty}\log\Phi\left(\frac{\sum_{k=2}^{t}\frac{k}{k+\sigma_{n}^{2}}(1-\Delta)}{\sqrt{\sum_{k=2}^{t}(1+\frac{\sigma_{n}^{2}}{k+\sigma_{n}^{2}})}}\right)$
$\geqslant\sum_{t=1}^{\infty}\log\Phi\left(\frac{(1-\Delta)\left(t-\sigma_{n}^{2}\ln{(t+\sigma_{n}^{2})}+2\sigma_{n}^{2}\ln{\sigma_{n}}-1/(\sigma_{n}^{2}+1)\right)}{\sqrt{t+\sigma_{n}^{2}\ln{(t+\sigma_{n}^{2})}-2\sigma_{n}^{2}\ln{\sigma_{n}}-(\sigma_{n}^{2}+2)/(\sigma_{n}^{2}+1)}}\right)$
[…]
$f_{t}(\Delta,\sigma_{n})=\frac{(1-\Delta)\left(t-\sigma_{n}^{2}\ln{(t+\sigma_{n}^{2})}+2\sigma_{n}^{2}\ln{\sigma_{n}}-1/(\sigma_{n}^{2}+1)\right)}{\sqrt{t+\sigma_{n}^{2}\ln{(t+\sigma_{n}^{2})}-2\sigma_{n}^{2}\ln{\sigma_{n}}-(\sigma_{n}^{2}+2)/(\sigma_{n}^{2}+1)}},$ (20)
Referring to the lower bound of the standard Gaussian distribution, we can continue to derive:
$\lim_{t\to\infty}\log\mathbb{P}(\Omega_{t})$ $\geqslant\sum_{t=1}^{\infty}\log\Phi(f_{t}(\Delta,\sigma_{n}))$ (21)
$\geqslant\sum_{t=1}^{\infty}\log\left(1-\frac{1}{\sqrt{2\pi}f_{t}(\Delta,\sigma_{n})e^{f_{t}^{2}(\Delta,\sigma_{n})/2}}\right)$
$\geqslant\sum_{t=1}^{\infty}\left(-\frac{1}{\sqrt{2\pi}f_{t}(\Delta,\sigma_{n})e^{f_{t}^{2}(\Delta,\sigma_{n})/2}}\right)$
>>$\infty$,> - ∞ ,
This shows that there exists a constant $c^{\prime}>0$ such that:
$\lim_{t\to\infty}\log\mathbb{P}(\Omega_{t})\geqslant\log c^{\prime}>-\infty$ (22)
In other words, we have:
$\lim_{t\to\infty}\mathbb{P}(\Omega_{t})\geqslant c>0,$ (23)
where $c=e^{c^{\prime}}$. Moreover, $\mathbb{P}(\Omega_{t})\geqslant\lim\limits_{t\to\infty}\mathbb{P}(\Omega_{t})\geqslant c$ for a finite sequence. Specifically, let $\Delta=0.1$ and $\sigma_{n}^{2}=0.1$, we can get $c^{\prime}=-0.62$, and $c=0.54$. ∎
Moreover, the function $f_{t}(\Delta,\sigma_{n}^{2})$, combined with the derivation above, demonstrates that the divergence probability $\mathbb{P}(\Omega_{t})$ decreases as $\Delta$ and $\sigma_{n}^{2}$ increase, which is consistent with our proposition (Proposition E.3) and the experiments.
The above argument is based on the frequentist setting where the underlying instance is fixed, and the agent does not access the right noise likelihood function of the environment. We conjecture that under the Bayesian setting where the prior and likelihood in the game environment are available to the agent, with the exact Bayes posterior, the TS-RM still suffers linear Bayesian adversarial regret.
E.3 Why Optimistic variant of TS can converge?
Assume that Alice chooses the wrong action until time $t$. By Prop E.3, the TS-RM algorithm will continue to choose the wrong action with a constant probability $c$. Different from TS-RM, we will prove that even if Alice chooses the wrong action until time $t$, OTS-RM will eventually yield a sub-linear regret with high probability.
Proof.
Unlike TR-RM, the OTS-RM algorithm uses $M$ samples for optimistic sampling in each round. Under the assumption that Alice takes the wrong action until time $t$, we have:
$\begin{cases}x_{ti}&\sim\mathcal{N}_{i}(0,1),\quad i=1,\ldots,M\\ y_{ti}&\sim\mathcal{N}_{i}\left(\frac{t}{t+\sigma_{n}^{2}}(1-\Delta),\frac{\sigma_{n}^{2}}{\sigma_{n}^{2}+t}\right),\quad i=1,\ldots,M\end{cases}$ (24)
Let $\tilde{R}_{t1}=\max\limits_{i}x_{ti}$ and $\tilde{R}_{t2}=\max\limits_{i}y_{ti}$, which are just $\tilde{R}_{t}(1nd)$ and $\tilde{R}_{t}(2nd)$ mentioned above, respectively.
The proof will first show that $\tilde{R}_{t1}\geqslant\tilde{R}_{t2}$ with high probability. As a result, $\Re_{t1}$ will decrease, and eventually Alice’s strategy will change from a pure strategy $[0,1]$ to a mixed strategy, indicating a decay of $\mathbb{P}(\Omega_{t})$ over time $t$.
According to the anti-concentration property in Lemma F.11, we have:
$\mathbb{P}(\max_{i}y_{ti}\leqslant\frac{t}{t+\sigma_{n}^{2}}(1-\Delta)+\sqrt{\frac{2\sigma_{n}^{2}\log(M/\delta_{1})}{t+\sigma_{n}^{2}}})\geqslant 1-\delta_{1}$ (25)
Thus, the first step in the proof can be written as:
$\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$ $=\mathbb{P}(\max_{i}x_{i}\geqslant\max_{i}y_{i})$ (26)
$\geqslant(1-\delta_{1})\mathbb{P}(\max_{i}x_{i}\geqslant\max_{i}y_{i}\mid\epsilon)$
$\geqslant(1-\delta_{1})\mathbb{P}(\max_{i}x_{ti}\geqslant\frac{t}{t+\sigma_{n}^{2}}(1-\Delta)+\sqrt{\frac{2\sigma_{n}^{2}\log(M/\delta_{1})}{t+\sigma_{n}^{2}}}\mid\epsilon)$
$\geqslant 1-\delta_{1}-(1-\delta_{1})\Phi^{M}\left(\frac{t}{t+\sigma_{n}^{2}}(1-\Delta)+\sqrt{\frac{2\sigma_{n}^{2}\log(M/\delta_{1})}{t+\sigma_{n}^{2}}}\right)$
Here, $\epsilon$ represents the event where the anti-concentration property occurs.
Let $f(t,\Delta,\sigma_{n})=\frac{t}{t+\sigma_{n}^{2}}(1-\Delta)+\sqrt{\frac{2\sigma_{n}^{2}\log(M/\delta_{1})}{t+\sigma_{n}^{2}}}$. Then, we have:
$\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$ $\geqslant 1-\delta_{1}-(1-\delta_{1})\Phi^{M}(f)$ (27)
$\geqslant 1-\delta_{1}-(1-\delta_{1})\left(1-\frac{f}{\sqrt{2\pi}(f^{2}+1)e^{f^{2}/2}}\right)^{M}$
$\geqslant 1-\delta_{1}-(1-\delta_{1})\exp{\left(\frac{-Mf}{\sqrt{2\pi}(f^{2}+1)e^{f^{2}/2}}\right)}$
where $\Phi(x)\leqslant 1-\frac{x}{\sqrt{2\pi}(x^{2}+1)e^{x^{2}/2}}$ (Gordon, 1941) and $(1-x)^{M}\leqslant e^{-Mx}$. ∎
To obtain further insights into the relationship between $M$ and $t$, we have depicted a figure in Figure 8 that corresponds to the inequality in Equation 27. Based on the analysis, we draw the following conclusions:
  • When $\Delta,\sigma_{n}$ and $t$ are fixed, $\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$ increases with $M$. Additionally, the value of $t$ has a significant influence on $\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$.
  • When $\Delta,\sigma_{n}$ and $M$ are fixed, the probability $\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$ increases with time $t$. Moreover, $\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$ will quickly reach a region close to the maximum in just a few rounds.
  • When $\sigma_{n},M$ and $t$ are fixed, $\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$ initially increases with $\Delta$ and subsequently decreases with $\Delta$.
    Figure 8: Relationship between $t$ and $M$ (fix $\Delta$ and $\sigma_{n}$)
These findings provide valuable insights into the behavior of the OTS-RM algorithm and support our claim that $\mathbb{P}(\tilde{R}_{t1}\geqslant\tilde{R}_{t2})$ increases as $M$ grows. This increasing probability implies that Alice’s strategy will transition from a pure strategy $[0,1]$ to a mixed strategy, indicating a decay of $\mathbb{P}(\Omega_{t})$ over time $t$. Consequently, the algorithm achieves sub-linear regret with high probability.
Appendix F Technical details in Section 4
Our analysis road map is as follows. First, in Section F.1, as described in Proposition 3.1, we derive a general regret decomposition in given any imagined reward sequence $\{\tilde{R}_{t}:t\in\mathbb{Z}_{++}\}$, where each $\tilde{R}_{t+1}$ is constructed using history information $H_{t}$ with algorithmic randomness. Then, to further upper bound the regret, we introduce the generic upper confidence bound (UCB) sequence and lower confidence bound (LCB) sequence as in Definition F.4. With these sequences, we have a general regret bound in Proposition F.8 with generic UCB and LCB sequences. Next, as described in Remark F.9, we specify the so-called information-theoretic confidence bound in Section F.2 and show that the imagined reward sequence has good properties when compared with specified sequences in Section F.3, finally yielding the information-theoretic regret upper bound in Section F.4.
F.1 General regret bound
Proposition F.1 (Restate regret decomposition in Proposition 3.1).
For any $a\in\mathcal{A}$, the one-step regret can be decomposed by
$\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right]=\mathbb{E}\left[f_{\theta}(a,B_{t})-f_{\theta}(A_{t},B_{t})\>|\>\theta\right]$
$=\underbrace{\mathbb{E}\left[\tilde{R}_{t+1}(a)-\tilde{R}_{t+1}(A_{t})\>|\>\theta\right]}_{(I)}+\underbrace{\mathbb{E}\left[f_{\theta}({a,B_{t}})-\tilde{R}_{t+1}(a)\>|\>\theta\right]}_{(II)}+\underbrace{\mathbb{E}\left[\tilde{R}_{t+1}(A_{t})-f_{\theta}(A_{t},B_{t})\>|\>\theta\right]}_{(III)}$
Proof.
Since
$\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right]=\mathbb{E}\left[f_{\theta}(a,B_{t})-f_{\theta}(A_{t},B_{t})\>|\>\theta\right]$
We have introduced the imagined time-varying sequence $\tilde{R}:=\{\tilde{R}_{t}:t\in\mathbb{Z}_{++}\}$, where each $\tilde{R}_{t}$ is constructed using history information $H_{t}$ and takes value in $[0,1]^{\mathcal{A}}$.
$f_{\theta}(a,B_{t})-f_{\theta}(A_{t},B_{t})=\underbrace{\tilde{R}_{t+1}(a)-\tilde{R}_{t+1}(A_{t})}_{(I)\leavevmode\nobreak\ \operatorname{adv}_{t+1}(a)}+\underbrace{f_{\theta}(a,B_{t})-\tilde{R}_{t+1}(a)}_{(II)\leavevmode\nobreak\ \operatorname{pess}_{t+1}(a)}+\underbrace{\tilde{R}_{t+1}(A_{t})-f_{\theta}(A_{t},B_{t})}_{(III)\operatorname{est}_{t+1}}$ (28)
Reduction to full-information adversarial regret.
Any algorithm $\pi=\pi^{\operatorname{adv-est}}$ constructs the imagined reward sequence $\tilde{R}$ and use $\operatorname{adv}$ algorithm for no-regret update will lead to,
$\sum_{t=0}^{T-1}\mathbb{E}\left[f_{\theta}(a,B_{t})-f_{\theta}(A_{t},B_{t})\>|\>\theta\right]$
$\leqslant\Re_{\operatorname{full}}(a;T,\operatorname{adv},\tilde{R})+\sum_{t=0}^{T-1}\mathbb{E}\left[\operatorname{pess}_{t+1}(a)\mid\theta\right]+\sum_{t=0}^{T-1}\mathbb{E}\left[\operatorname{est}_{t+1}\>|\>\theta\right]$ (29)
where
$\Re_{\operatorname{full}}(a;T,\operatorname{adv},\tilde{R})$ $=\mathbb{E}\left[\sum_{t=0}^{T-1}\tilde{R}_{t+1}(a)-\tilde{R}_{t+1}(A_{t})\right].$
Recall the Bayesian adversarial regret is
$\Re(T,\pi^{\operatorname{alg}},\pi^{B})=\mathbb{E}\left[\Re(T,\pi^{\operatorname{alg}},\pi^{B},\theta)\right]$
where the expectation is taken over the prior distribution of $\theta$. From Section F.1, we have
$\Re(T,\pi^{\operatorname{alg}},\pi^{B})$ $=\mathbb{E}\left[\max_{a\in\mathcal{A}}\sum_{t=0}^{T-1}\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\mid\theta\right]\right]$
$\leqslant\max_{a\in\mathcal{A}}\Re_{\operatorname{full}}(a;T,\operatorname{adv},\tilde{R})+\mathbb{E}\left[\max_{a\in\mathcal{A}}\sum_{t=0}^{T-1}\mathbb{E}\left[\operatorname{pess}_{t+1}(a)\mid\theta\right]\right]+\sum_{t=0}^{T-1}\mathbb{E}\left[\operatorname{est}_{t+1}\right]$
$=\Re_{\operatorname{full}}(T,\operatorname{adv},\tilde{R})+\mathbb{E}\left[\max_{a\in\mathcal{A}}\sum_{t=0}^{T-1}\mathbb{E}\left[\operatorname{pess}_{t+1}(a)\mid\theta\right]\right]+\sum_{t=0}^{T-1}\mathbb{E}\left[\operatorname{est}_{t+1}\right]$ (30)
Remark F.2.
If each imagined reward $\tilde{R}_{t}$ in the sequence $\tilde{R}$ takes bounded value in $[-c,c]^{\mathcal{A}}$, any suitable full information adversarial algorithm $\operatorname{adv}$ will give satisfied bound for $\Re_{\operatorname{full}}(a;T,\operatorname{adv},\tilde{R})$ for any $a\in\mathcal{A}$. Specifically, Hedge suffers $\Re_{\operatorname{adv}}(a;T,\text{Hedge},\tilde{R})=\mathcal{O}(2c\sqrt{T\log\mathcal{A}})$ and RM suffers $\Re_{\operatorname{full}}(a;T,\text{RM},\tilde{R})=\mathcal{O}(2c\sqrt{T\mathcal{A}})$.
Next, we focus on the pessimism term and estimation term. We now focus on the case generalized from the optimistic Thompson sampling.
Assumption F.3 (Restriction on the imagined reward sequence $\tilde{R}$).
In the following context, the imagined reward sequence $\tilde{R}=(\tilde{R}_{1},\ldots,\tilde{R}_{t+1},\ldots)$ satisfies that (1) $\tilde{R}_{t+1}(a)\in[0,C]$ for all $a\in\mathcal{A}$ and (2) $\tilde{R}_{t+1}(a)$ is random only through its dependence on $Z_{t+1}$ given the history $H_{t},B_{t}$ for all $a\in\mathcal{A}$. To clarify, $\tilde{R}_{t+1}(a)$ has no dependence on $A_{t}$ and $\tilde{R}_{t+1,A_{t},B_{t}}$.
Definition F.4 (UCB and LCB sequence).
UCB sequence $U=(U_{t}\>|\>t\in\mathbb{N})$ LCB sequence $L=(L_{t}\>|\>t\in\mathbb{N})$ are two sequences of functions where each $0\leqslant L_{t}\leqslant U_{t}\leqslant C$ are both deterministic given history $H_{t}$.
Definition F.5 (Optimistic Event).
For any imagined reward sequence $\tilde{R}_{t+1}$ in F.3 and any upper confidence sequence $U$ in Definition F.4, we define the event
$\mathcal{E}_{t}(\tilde{R},U,B_{t}):=\{\tilde{R}_{t+1}(a)\geqslant U_{t}(a,B_{t}),\forall a\in\mathcal{A}\}.$
Fact F.1.
The event $\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})$ is random only through its dependence on $Z_{t+1}$ given $H_{t},B_{t}$. The pessimism term can be decomposed according to the event $\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})$
$\tilde{R}_{t+1}(a)-f_{\theta}(a,B_{t})=\mathds{1}_{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})}(\tilde{R}_{t+1}(a)-f_{\theta}(a,B_{t}))+(1-\mathds{1}_{{\mathcal{E}}^{o}_{t}(\tilde{R},U,B_{t})})(\tilde{R}_{t+1}(a)-f_{\theta}(a,B_{t}))$
Consider the case where $f_{\theta}$ takes values in $[0,C]$ and $\tilde{R}_{t+1}\in[0,C]$ by F.3, for all $a\in\mathcal{A}$,
$f_{\theta}(a,B_{t})-\tilde{R}_{t+1}(a)\leqslant\mathds{1}_{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})}(f_{\theta}(a,B_{t})-U_{t}(a,B_{t}))+C(1-\mathds{1}_{{\mathcal{E}}^{o}_{t}(\tilde{R},U,B_{t})}).$ (31)
Definition F.6 (Concentration Event).
For any imagined reward sequence $\tilde{R}_{t+1}$ in F.3 and any upper confidence sequence $U^{\prime}$ in Definition F.4, we define the event
$\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t}):=\{\tilde{R}_{t+1}(A_{t})\leqslant U^{\prime}_{t}(A_{t},B_{t})\}.$
Fact F.2.
Consider the case where $f_{\theta}$ takes values in $[0,C]$ and $\tilde{R}_{t+1}\in[0,C]$ by F.3, the event $\mathcal{E}_{t}(\tilde{R},U^{\prime},A_{t},B_{t})$ is random only through its dependence on $Z_{t+1}$ given $H_{t},A_{t},B_{t}$. The estimation term then becomes,
$\tilde{R}_{t+1}(A_{t})-f_{\theta}(A_{t},B_{t})\leqslant(\mathds{1}_{\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t})})(U^{\prime}_{t}(A_{t},B_{t})-f_{\theta}(A_{t},B_{t}))+C(1-\mathds{1}_{\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t})}).$ (32)
Definition F.7 (Confidence event).
Define the confidence event at round $t$ as
$\mathcal{E}_{t}(f_{\theta},B_{t}):=\{\forall a\in\mathcal{A},f_{\theta}(a,B_{t})\in[L_{t}(a,B_{t}),U_{t}(a,B_{t})]\}.$
Fact F.3.
The event $\mathcal{E}^{c}_{t}(f_{\theta},B_{t})$ is deterministic conditioned on history $H_{t}$, $B_{t}$ and $\theta$.
Based on the definition of the confidence event, the pessimism term from Equation 31 becomes
$f_{\theta}(a,B_{t})-\tilde{R}_{t+1}(a)\leqslant C(1-\mathds{1}_{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})\cap\mathcal{E}^{c}_{t}(f_{\theta},B_{t})}),\quad\forall a\in\mathcal{A}.$
The estimation term from Equation 32 becomes
$\tilde{R}_{t+1}(A_{t})-f_{\theta}(A_{t},B_{t})$ $\leqslant\mathds{1}_{\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t})\cap\mathcal{E}^{c}_{t}(f_{\theta},B_{t})}(U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t}))$
$\quad+C\left(1-\mathds{1}_{\mathcal{E}^{c}_{t}(\tilde{R},U,A_{t},B_{t})}\mathds{1}_{\mathcal{E}^{c}_{t}(f_{\theta},B_{t})}\right)$
Denote the complement event of $\mathcal{E}$ as $\neg{\mathcal{E}}$.
$\mathbb{E}\left[\max_{a\in\mathcal{A}}\sum_{t=0}^{T-1}\mathbb{E}\left[\operatorname{pess}_{t+1}(a)\mid\theta\right]\right]$ $\leqslant\mathbb{E}\left[\max_{a\in\mathcal{A}}\sum_{t=0}^{T-1}\mathbb{E}\left[C(1-\mathds{1}_{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})\cap\mathcal{E}^{c}_{t}(f_{\theta},B_{t})})\mid\theta\right]\right]$
$=\mathbb{E}\left[\sum_{t=0}^{T-1}\mathbb{E}\left[C(1-\mathds{1}_{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})\cap\mathcal{E}^{c}_{t}(f_{\theta},B_{t})})\right]\right]$
$=C\sum_{t=0}^{T-1}\mathbb{P}\left(\neg{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})}\cup\neg{\mathcal{E}^{c}_{t}(f_{\theta},B_{t})}\right)$ (33)
and by assuming $U^{\prime}_{t}\geqslant L_{t}$
$\mathbb{E}\left[\sum_{t=0}^{T-1}\operatorname{est}_{t+1}\right]$ $\leqslant\sum_{t=0}^{T-1}\mathbb{E}\left[(U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t}))\right]$
$+C\sum_{t=0}^{T}\mathbb{P}(\neg\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t})\cup\neg\mathcal{E}^{c}_{t}(f_{\theta},B_{t}))$ (34)
Proposition F.8 (General Regret Bound with confidence sequence.).
Given sequences $U^{\prime}\geqslant U\geqslant L$, we upper bound the Bayesian adversarial regret with
$\Re(T,\pi^{\operatorname{alg}},\pi^{B})$ $\leqslant\Re_{\operatorname{full}}(T,\operatorname{adv},\tilde{R})+\sum_{t=0}^{T-1}\mathbb{E}\left[(U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t}))\right]$
$\quad+C\sum_{t=0}^{T-1}\mathbb{P}(\neg\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t}))+2\mathbb{P}(\neg\mathcal{E}^{c}_{t}(f_{\theta},B_{t}))+\mathbb{P}\left(\neg{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})}\right)$ (35)
Proof.
This is a direct consequence of Equations 30, F.1 and F.1. ∎
Remark F.9.
In the following sections, we will discuss the specific choice of the UCB and LCB sequence $U^{\prime},U,L$. In Section F.2, we will show that with information-theoretic confidence bound, the probability ($\mathbb{P}(\neg\mathcal{E}^{c}_{t}(f_{\theta},B_{t}))$ of the function $f_{\theta}$ not covered in the confidence region is small. In Section F.3, we will show the imagined reward sequence constructed by OTS and UCB will lead to small will stay within the sequence $U^{\prime}$ and $U$ with high probability, i.e., the probability ($\mathbb{P}\left(\neg{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})}\right)$ and ($\mathbb{P}(\neg\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t}))$ is small enough. In Section F.4, we show that $\sum_{t=0}^{T-1}\mathbb{E}\left[(U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t}))\right]$ can be bounded by the mutual information $I(\theta;H_{T})$ with the information-theoretic confidence bound defined in Section F.2.
F.2 Information-theoretic confidence bound
Fact F.4 (Chernoff bound).
Suppose $X$ is normal distributed $N(\mu,\sigma^{2})$, the optimized chernoff bound for $X$ is
$\mathbb{P}(X-\mu\geqslant c)$ $\leqslant\min_{t>0}\frac{\exp(\sigma^{2}t^{2}/2)}{\exp(tc)}=\exp(-c^{2}/\sigma^{2})$
Lemma F.10.
Conditioned on $H_{t}$ and $B_{t}$, Define the set $\mathcal{F}_{t}$ as
$\mathcal{F}_{t}:=\left\{f_{\theta}:\left|f_{\theta}(a,B_{t})-\mu_{t}(a,B_{t})\right|\leqslant\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,B_{t}),\forall a\in\mathcal{A}\right\}$
Then,
$\mathbb{P}(f_{\theta}\in\mathcal{F}_{t})\geqslant 1-2\mathcal{A}\exp(-\beta^{\prime}_{t}/2).$
Proof.
Since $f_{\theta}(a,b)\mid H_{t}$ is distributed as $N(\mu_{t}(a,b),\sigma_{t}(a,b))$, by F.4,
$\mathbb{P}(\left|f_{\theta}(a,b)-{\mu_{t}(a,b)}\right|\geqslant\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,b)\mid H_{t})$ $\leqslant 2\exp\left(-\frac{\beta^{\prime}_{t}}{2}\right)$ (36)
By union bound, we have
$\mathbb{P}(\left|f_{\theta}(a,b)-{\mu_{t}(a,b)}\right|\geqslant\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,b),\forall a\in\mathcal{A}\mid H_{t})\leqslant 2\mathcal{A}\exp(-\beta^{\prime}_{t}/2),$
for any fixed $b\in\mathcal{B}$, We observe that conditioned on $H_{t}$, the opponent’s action $B_{t}$ is independent of $f_{\theta}(a,b)$ for all $(a,b)\in\mathcal{A}\times\mathcal{B}$. Therefore, we further derive
$\mathbb{P}(\left|f_{\theta}(a,B_{t})-{\mu_{t}(a,B_{t})}\right|\geqslant\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,b),\forall a\in\mathcal{A}\mid H_{t})\leqslant 2\mathcal{A}\exp(-\beta^{\prime}_{t}/2).$
Taking expectations on both sides, we prove the lemma. ∎
Let the UCB and LCB sequences be
$U=(\mu_{t}(a,b)+\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,b):t\in\mathbb{N}),\quad L=(\mu_{t}(a,b)-\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,b):t\in\mathbb{N}).$
Let
$\beta^{\prime}_{t}=2\log\mathcal{A}\sqrt{t}.$
By Lemma F.10, we can see the probability introduced in Definition F.7 is ($\mathbb{P}(\neg\mathcal{E}^{c}_{t}(f_{\theta},B_{t}))\leqslant 2\mathcal{A}\exp(-\beta_{t}^{\prime}/2)=2/\sqrt{t}$.
Relate $\sigma_{t}$ to information-theoretic quantity.
Fact F.5 (Mutual information of Gaussian distribution).
If $f_{\theta}(a)\sim N(\mu(a),\sigma(a))$ and $o=f_{\theta}(a)+w$ with fixed $a$ and $w\sim N\left(0,\sigma_{w}^{2}\right)$, then
$I(\theta;o)=\frac{1}{2}\log\left(1+{\sigma_{w}^{-2}\sigma(a)}\right)$
Proof.
$I(\theta;o)=H(o)-H(o\mid\theta)=\frac{1}{2}\log 2\pi e(\sigma(a)^{2}+\sigma_{w}^{2})-\frac{1}{2}\log 2\pi e\sigma_{w}^{2}=\frac{1}{2}\log(1+\sigma_{w}^{-2}\sigma(a))$
For an fixed action pair $(a,b)$, by the F.5,
$I_{t}\left(\theta;R_{t+1,A_{t},B_{t}}\mid A_{t}=a,B_{t}=b\right)=\frac{1}{2}\log\left(1+\frac{\sigma^{2}_{t}(a,b)}{\sigma_{w}^{2}}\right)$
Let the width function $w_{t}$ be
$w_{t}(a,b)=\sqrt{\beta_{t}I_{t}\left(\theta;R_{t+1,A_{t},B_{t}}\mid A_{t}=a,B_{t}=b\right)}\quad\text{ with }\quad\beta_{t}=\frac{2\beta^{\prime}_{t}}{\log(1+\sigma_{w}^{-2})}.$
Expanding $w_{t}(a,b)$ leads to
$w_{t}(a,b)^{2}=\beta_{t}I_{t}(\theta;R_{t+1,A_{t},B_{t}}\>|\>A_{t}=a,B_{t}=b)=\frac{\beta^{\prime}_{t}\log(1+\sigma_{w}^{-2}\sigma^{2}_{t}(a,b))}{\log(1+\sigma_{w}^{-2})}\geqslant\beta^{\prime}_{t}\sigma^{2}_{t}(a,b).$ (37)
The last inequality follows from the fact that $\frac{x}{\log(1+x)}$ monotonically increases for $x>0$, and that $\sigma^{2}_{t}(a,b)\leqslant k((a,b),(a,b))\leqslant 1$, leading to the result $\sigma_{t}^{2}(a,b)\leqslant\frac{1}{\log(1+\sigma_{w}^{-2})}\log(1+\sigma_{w}^{-2}\sigma^{2}_{t}(a,b)).$
F.3 Anti-concentration behavior of optimistic Thompson sampling
Let the UCB sequences be
$U=(\mu_{t}(a,b)+\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,b):t\in\mathbb{N}),\quad U^{\prime}=(\mu_{t}(a,b)+\sqrt{2\log(M\sqrt{t})}\sigma_{t}(a,b):t\in\mathbb{N}).$
In this section, we show the probability ($\mathbb{P}\left(\neg{\mathcal{E}^{o}_{t}(\tilde{R},U,B_{t})}\right)$ and ($\mathbb{P}(\neg\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t}))$ is small enough. The following two lemmas are important.
Lemma F.11.
Consider a normal distribution $N\left(0,\sigma^{2}\right)$ where $\sigma$ is a scalar. Let $\eta_{1},\eta_{2},\ldots,\eta_{M}$ be $M$ independent samples from the distribution. For any $w\in\mathbb{R}_{+}$,
$\mathbb{P}\left(\max_{j\in[M]}\eta_{j}\geqslant w\right)=1-\left[\Phi\left(\frac{w}{\sigma}\right)\right]^{M}$
Proof.
By the fact of Normal distribution, $\mathbb{P}\left(\eta_{j}\leqslant w\right)=\mathbb{P}({\eta_{j}}/{\sigma}\leqslant{w}/{\sigma})=\Phi({w}/{\sigma}).$ We have
$\mathbb{P}(\max_{j\in[M]}\eta_{j}\geqslant w)=1-\mathbb{P}(\max_{j\in[M]}\eta_{j}\leqslant w)=1-\mathbb{P}(\forall j\in[M],\eta_{j}\leqslant w)=1-\left[\Phi\left(\frac{w}{\sigma}\right)\right]^{M}.$
Proposition F.12.
Let the UCB sequence be
$U=(\mu_{t}(a,b)+\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,b):t\in\mathbb{N}).$
Let
$M=\frac{\log(\sqrt{t})}{\log\frac{1}{\Phi(\sqrt{\beta^{\prime}_{t}})}}$
and we have,
($\mathbb{P}(\neg\mathcal{E}_{t}(\tilde{R},U,B_{t}))\leqslant\frac{1}{\sqrt{t}}.$
The summation of failure probability is
$\sum_{t=1}^{T}\frac{1}{\sqrt{t}}=2\sum_{t=1}^{T}\frac{t-(t-1)}{\sqrt{t}+\sqrt{t}}\leqslant 2\sum_{t=1}^{T}\frac{t-(t-1)}{\sqrt{t}+\sqrt{t-1}}=2\sum_{t=1}^{T}(\sqrt{t}-\sqrt{t-1})=2\sqrt{T}$
Proof.
Recall that the optimistic Thompson sampling estimator $\tilde{R}$ is generated by
$\tilde{f}^{\operatorname{OTS}}_{t+1}(a,B_{t}):=(\max_{j\in[M]}z_{t+1}^{j})\cdot\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,B_{t})+\mu_{t}(a,B_{t})\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \tilde{R}_{t+1}(a)=\operatorname{clip}_{[-c,c]}\left(\tilde{f}_{t+1}^{\operatorname{OTS}}(a,B_{t})\right),\forall a\in\mathcal{A}.$
$\mathbb{P}(\mathcal{E}_{t}(\tilde{R},U,B_{t})\mid H_{t},B_{t})$ $=\mathbb{P}(\tilde{R}_{t+1}(a)\geqslant U_{t}(a,B_{t}),\forall a\in\mathcal{A}\mid H_{t},B_{t})$
$\stackrel{{\scriptstyle(i)}}{{\geqslant}}\mathbb{P}(\tilde{R}_{t+1}(a)\geqslant\min\{\mu_{t}(a,B_{t})+\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,B_{t}),c\},\forall a\in\mathcal{A}\>|\>H_{t},B_{t})$
$\stackrel{{\scriptstyle(ii)}}{{\geqslant}}\mathbb{P}\left(\left(\max_{j\in[M]}z^{j}_{t+1}\right)\sigma_{t}(a,B_{t})\geqslant\sqrt{\beta^{\prime}_{t}}\sigma_{t}(a,B_{t}),\forall a\in\mathcal{A}\>|\>H_{t},B_{t}\right)$
$=\mathbb{P}\left(\max_{j\in[M]}z^{j}_{t+1}\geqslant\sqrt{\beta^{\prime}_{t}}\right)$
=$\stackrel{{\scriptstyle(iii)}}{{=}}1-\Phi\left(\sqrt{\beta_{t}^{\prime}}\right)^{M}.$ (38)
where $(i)$ is by the definition of $\mathcal{E}(c)$ and $\mathcal{F}_{t}$ and the fact $B_{t}$ is conditionally independent of $f_{\theta}$ given $H_{t}$; $(ii)$ is due to $\operatorname{clip}_{[-c,c]}(x)\geqslant\min(x,c)$ and the function $min(x,c)$ is non-decreasing function on $x$; $(iii)$ is by the fact $\{z^{j}_{t+1}\}_{j\in[M]}$ is independent of $H_{t+1}=(H_{t},A_{t},B_{t},R_{t+1,A_{t},B_{t}})$, the fact that $\sigma_{t}(a,b)$ and $w_{t}(a,b)$ are deterministic given $H_{t}$ and Lemma F.11. Solve $\Phi(\beta^{\prime}_{t})^{M}=1/\sqrt{t}$, we have $M=\frac{\log(\sqrt{t})}{\log\frac{1}{\Phi(\sqrt{\beta^{\prime}_{t}})}}.$
Lemma F.13 (Anti-concentration property of maximum of Gaussian R.V.).
Consider a normal distribution $N\left(0,\sigma^{2}\right)$ where $\sigma$ is a scalar. Let $\eta_{1},\eta_{2},\ldots,\eta_{M}$ be $M$ independent samples from the distribution. Then for any $\delta>0$,
$\mathbb{P}\left(\max_{j\in[M]}\eta_{j}\leqslant\sqrt{2\sigma^{2}\log(M/\delta)}\right)\geqslant 1-\delta.$
Proposition F.14.
Let the UCB sequences be
$U^{\prime}=(\mu_{t}(a,b)+\sqrt{2\log(M\sqrt{t})}\sigma_{t}(a,b):t\in\mathbb{N}).$
Then
($\mathbb{P}(\neg\mathcal{E}^{c}_{t}(\tilde{R},U^{\prime},A_{t},B_{t}))\leqslant\frac{1}{\sqrt{t}}.$
Proof.
This is a direct consequence of Lemma F.13 under conditional probability given $H_{t},A_{t},B_{t}$ with setting $\delta=1/\sqrt{t}$. ∎
F.4 Bounding Estimation Regret via Information-theoretic quantity
To characterize the property of game environments and how much information algorithm can acquire about the environment at each round, we define the information ratio of algorithm $\pi=(\pi_{t})_{t\in\mathbb{N}}$,
From Equation 37, we have
$\beta_{t}I_{t}(\theta;R_{t+1,A_{t},B_{t}}\>|\>A_{t}=a,B_{t}=b)\geqslant\beta^{\prime}_{t}\sigma^{2}_{t}(a,b).$ (39)
where
$\beta_{t}=\frac{2\beta^{\prime}_{t}}{\log(1+\sigma_{w}^{-2})}.$
Recall that
$U^{\prime}=(\mu_{t}(a,b)+\sqrt{2\log(M\sqrt{t})}\sigma_{t}(a,b):t\in\mathbb{N}).$
Then,
$(U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t}))$ $\leqslant\left(\frac{\sqrt{2\log(M\sqrt{t})}}{\sqrt{\beta^{\prime}_{t}}}+1\right)\sqrt{\beta^{\prime}_{t}}\sigma_{t}(A_{t},B_{t})$
$\leqslant\left(\frac{\sqrt{2\log(M\sqrt{t})}}{\sqrt{\beta^{\prime}_{t}}}+1\right)\sqrt{\beta_{t}I_{t}(\theta;R_{t+1,A_{t},B_{t}}\>|\>A_{t},B_{t})}$
$=\left(\frac{\sqrt{2\log(M\sqrt{t})}}{\sqrt{\beta^{\prime}_{t}}}+1\right)\sqrt{\beta_{t}I_{t}(\theta;A_{t},B_{t},R_{t+1,A_{t},B_{t}})}$
where $\frac{\sqrt{2\log M\sqrt{t}}}{\sqrt{\beta^{\prime}_{t}}}=\sqrt{\frac{2\log M\sqrt{t}}{\log\mathcal{A}\sqrt{t}}}$. Then immediately from Proposition F.8 and Cauchy–Bunyakovsky–Schwarz inequality,
$\mathbb{E}\left[\sum_{t=0}^{T-1}U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t})\right]$ $\leqslant\mathbb{E}\left[\sum_{t=0}^{T-1}\sqrt{\left(\sqrt{2\log(M\sqrt{t})/\beta^{\prime}_{t}}+1\right)^{2}\beta_{t}I_{t}(\theta;A_{t},B_{t},R_{t+1,A_{t},B_{t}})}\right]$
$\leqslant\sqrt{\mathbb{E}\left[\sum_{t=0}^{T-1}\left(\sqrt{2\log(M\sqrt{t})/\beta^{\prime}_{t}}+1\right)^{2}\beta_{t}\right]}\sqrt{\mathbb{E}\left[\sum_{t=0}^{T-1}I_{t}(\theta;A_{t},B_{t},R_{t+1,A_{t},B_{t}})\right]}$
Let $Z_{t}=(A_{t},B_{t},R_{t+1,A_{t},B_{t}}))$, then
$\mathbb{E}\left[\sum_{t=0}^{T-1}I_{t}(\theta;A_{t},B_{t},R_{t+1,A_{t},B_{t}})\right]$ $=\mathbb{E}\left[\sum_{t}I_{t}(\theta;Z_{t})\right]=\sum_{t=0}^{T-1}I(\theta;Z_{t}\>|\>Z_{0},\ldots,Z_{t-1})$
$=I(\theta;Z_{0},\ldots,Z_{T-1})=I(\theta;H_{T})$
Then,
$\mathbb{E}\left[\sum_{t=0}^{T-1}U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t})\right]$ $\leqslant\sqrt{T\left(8\log(M\sqrt{T})/\log(1+\sigma_{w}^{-2})+2\beta_{T}\right)I(\theta;H_{T})}$
Plugin the choice of $M=\frac{\log(\mathcal{A}\sqrt{T})}{\log\frac{1}{\Phi(\beta^{\prime}_{t})}},\beta^{\prime}_{t}=2\log\mathcal{A}\sqrt{T}$ and $\beta_{t}=2\beta^{\prime}_{t}/\log(1+\sigma_{w}^{-2})$, then we obtain the final result,
$\log(M\sqrt{T})/\log(1+\sigma_{w}^{-2})+2\beta_{T}=\mathcal{O}(\log\mathcal{A}T+\log T+\log\log\mathcal{A}T).$
That is
$\mathbb{E}\left[\sum_{t=0}^{T-1}U^{\prime}_{t}(A_{t},B_{t})-L_{t}(A_{t},B_{t})\right]$ $\leqslant\mathcal{O}\left(\sqrt{T\log(\mathcal{A}T)I(\theta;H_{T})}\right).$ (40)
F.4.1 Bounds on the information gain $I(\theta;H_{T})$.
Remark F.15.
An important property of the Gaussian distribution is that the information gain does not depend on the observed rewards. This is because the posterior covariance of a multivariate Gaussian is a deterministic function of the points that were sampled. For this reason, this maximum information ratio $\gamma_{T}$ in Definition 4.3 is well defined. That is, $I(\theta;H_{T})=I(\theta;A_{0},B_{0},\ldots,A_{T-1},B_{T-1})\leqslant\gamma_{T}.$
We adopt the results from (Srinivas, Krause, Kakade, and Seeger, 2009) which gives the bounds of $\gamma_{T}$ for a range of commonly used covariance functions: finite dimensional linear, squared exponential and Matern kernels.
Example F.1 (Finite dimensional linear kernels).
Finite dimensional linear kernels have the form $k(x,x^{\prime})=x^{\top}x^{\prime}$. GPs with this kernel correspond to random linear functions $f(x)=\theta^{\top}x,\theta\sim N(0,\sigma_{0}I)$.
Example F.2 (Squared exponential kernel).
The Squared Exponential kernel is $k(x,x^{\prime})=\exp(-(2l^{2})^{-1}\left\|x-x^{\prime}\right\|^{2}s)$, $l$ is a lengthscale parameter. Sample functions are differentiable to any order almost surely.
Example F.3 (Matern kernel).
The Matern kernel is given by $k(x,x^{\prime})=(2^{1-\nu}/\Gamma(\nu))r^{\nu}B_{\nu}(r)$ and $r=(\sqrt{2\nu}/l)\left\|x-x^{\prime}\right\|$, where $\nu$ controls the smoothness of sample paths (the smaller, the rougher) and $B_{\nu}$ is a modified Bessel function. Note that as $\nu\rightarrow\infty$, appropriately rescaled Matern kernels converge to the Squared Exponential kernel.
Table 5: Maximum information gain $\gamma_{T}$.
Kernel Linear Squared exponential Materns ($\nu>1$)
$\gamma_{T}$ $\mathcal{O}(d\log T)$ $\mathcal{O}((\log T)^{d+1})$ $\mathcal{O}(T^{d(d+1)/(2\nu+d(d+1))}(\log T))$
Appendix G Experiments
G.1 Performance metric
Two performance metrics are utilized to evaluate the algorithms: average regret and KL divergence to Nash equilibrium.
Average regret:
The expected regret of player $i$ over $T$ time steps is defined as:
$\text{Regret}^{i}(T)=\frac{1}{T}\max_{a\in\Delta^{\mathcal{D}(\mathcal{A}^{i})}}\mathbb{E}\left[\sum_{t=1}^{T}\phi\left(a,x_{t}^{-i}\right)-\phi\left(x_{t}^{i},x_{t}^{-i}\right)\right],$ (41)
where $a$ and $x_{t}^{i}$ represent the strategy for player $i$ and $x_{t}^{-i}$ denotes the strategies of all other players. $\mathcal{D}(\mathcal{A}^{i})$ is the action set for player $i$. The expectation is taken over the randomness of the algorithm and the environment.
Duality Gap &\&& KL divergence to Nash:
The duality gap for a strategy pair $(x,y)$ is defined as
$\text{Gap}(x,y)=\max_{\left(x^{\prime},y^{\prime}\right)\in\Delta}\mathbb{E}\left[\phi\left(x^{\prime},y\right)-\phi\left(x,y^{\prime}\right)\right]$
The duality gap provides a measure of how close a solution pair is to a Nash equilibrium. If a solution pair $(x,y)$ has a duality gap of $\epsilon$, it is considered an $\epsilon$-Nash Equilibrium.
After $T$ iterations, the average-iterate strategies are defined as:
$\overline{x}_{T}=\frac{1}{T}\sum_{t=1}^{T}x_{t},\quad\overline{y}_{T}=\frac{1}{T}\sum_{t=1}^{T}y_{t}.$
The KL divergences $KL(\overline{x}_{T},x^{)}$ and $KL(\overline{y}_{T},y^{)}$ are also used as performance metrics for comparing the average-iterate solution pair $(\overline{x}_{T},\overline{y}_{T})$ with a Nash equilibrium pair $(x^{,}y^{)}$.
G.2 Different opponents in matrix game
Four types of opponents in random matrix games are introduced, and the performance of different algorithms against these opponents is compared.
Self-play opponent:
The opponent uses the same algorithm as the player.
Best-response opponent:
The strategy for a best-response opponent is defined as:
$y^{*}=\operatorname*{arg\,min}\limits_{y\in\Delta}y^{T}(Ax),$ (42)
which implies that the opponent knows matrix $A$ and the player’s strategy $x$ at each round.
Stationary opponent:
The stationary opponent always samples an action from a fixed strategy.
Non-stationary opponent:
A non-stationary opponent changes its strategy randomly every round.
G.3 Devices
Intel(R) Xeon(R) God 6230R CPU @ 2.10GHz and RTX A5000.
G.4 Additional results for random matrix games
This section presents additional evaluations of different algorithms on two-player zero-sum matrix games. The experiments consider payoff matrices where each entry is an i.i.d. random variable generated from the uniform distribution $[-1,1]$. Each player has $M$ actions, resulting in a squared payoff matrix of size $M$. The total number of rounds is set to $T=10^{7}$. In each round, the players receive noisy rewards ±$\pm\tilde{r}_{t}$, where $\tilde{r}t=A{ij}+\epsilon_{t}$, and $\epsilon_{t}\sim\mathcal{N}(0,0.1)$. The experiments investigate different matrix sizes, specifically $M=5,10,20,50,70,100$, and for each choice of $M$, $100$ independent simulation runs are conducted. The performance of the algorithms is averaged over these simulation runs.
Self-play opponent
First, we compare different algorithms under the self-play setting, where both players employ the same algorithm. Convergence curves of two performance metrics (see Section G.1) are shown in Figure 9, where each subplot (a)-(f) corresponds to a different matrix size $M$. The results indicate that algorithms exploiting the game structure outperform the two IWE baselines, particularly for smaller matrix sizes. As the matrix dimension increases, the performance gap between the proposed algorithms and the baselines diminishes. Among the algorithms, those based on the OTS method exhibit faster convergence than the ones based on UCB. Additionally, the average regret of RM decreases earlier than that of Hedge.
(a) $5\times 5$
(b) $10\times 10$
(c) $20\times 20$
(d) $50\times 50$
(e) $70\times 70$
(f) $100\times 100$
Figure 9: Self-play on different random matrix.
Best-response opponent
In this subsection, we introduce a best-response opponent (see Section G.2), while the player continues to use the various algorithms described above. Figure 10 presents the results for different matrix sizes. We observe that all the algorithms in our proposed framework outperform the IWE baselines. Once again, the OTS-based algorithms demonstrate a faster convergence behavior compared to UCB.
(a) $5\times 5$
(b) $10\times 10$
(c) $20\times 20$
(d) $50\times 50$
(e) $70\times 70$
(f) $100\times 100$
Figure 10: Best-response opponent on different random matrix.
Stationary opponent
Here, we consider a stationary opponent whose strategy remains fixed as a probability simplex over the action space, with values generated from a uniform distribution. The average regret in this scenario reflects the ability to exploit the opponent’s weakness. Convergence curves of the two performance metrics are compared in Figure 11. The results clearly demonstrate that the OTS estimator provides a significant advantage in exploiting this weak opponent, in comparison to the IWE-based estimators.
(a) $5\times 5$
(b) $10\times 10$
(c) $20\times 20$
(d) $50\times 50$
(e) $70\times 70$
(f) $100\times 100$
Figure 11: Stationary opponent on different random matrix.
Non-stationary opponent
In this subsection, we introduce a non-stationary opponent, requiring the player to develop a robust strategy against all possible opponent’s strategies. Specifically, a game matrix $A\in\mathbb{R}^{10\times 5}$ is generated, with each element sampled from $\mathcal{N}(0.5,2.0)$. The opponent’s actions are drawn from a fixed strategy that randomly changes every $50$ rounds. Each algorithm is evaluated over $1000$ rounds, and $100$ simulation runs are conducted. Figure 12 presents histograms of rewards over all rounds and simulation runs, while Table 6 summarizes the percentage of negative rewards and the mean reward values. The results show that OTS has a smaller percentage of negative rewards and achieves higher mean rewards compared to IWE, indicating its superior robustness.
Figure 12: Reward histograms of different algorithms against a non-stationary opponent.
return$<0$ mean return
IWE-Hedge 19.4% 1.24
IWE-RM 12.6% 1.50
OTS-Hedge 2.5% 1.55
OTS-RM 8.8% 1.55
Table 6: Returned rewards of IWE and OTS against a non-stationary opponent.
G.5 Sensitivity test for UCB-RM and OTS-RM
In this subsection, we have conducted new experiments involving the selection of different parameters within the UCB-Regret Matching$(\beta)$ and OTS-Regret $(M)$. These experiments encompass different matrix sizes (10(10( 10, $20$, and 50)50)50 ) under the self-play setting. The choices of $\beta$ are as follows, presented in descending order: $\sqrt{2\log\mathcal{A}\sqrt{T}},0.2\sqrt{2\log\mathcal{A}\mathcal{B}\sqrt{T}},0.05\sqrt{2\log\mathcal{A}\mathcal{B}\sqrt{T}}$, and0.$0.01\sqrt{2\log\mathcal{A}\mathcal{B}\sqrt{T}}$. Notably, $\sqrt{2\log\mathcal{A}\sqrt{T}}$ stems from our proof, while the subsequent selections are smaller values based on $\sqrt{2\log\mathcal{A}\mathcal{B}\sqrt{T}}$. The results demonstrate that decreasing the value of $\beta$ yields an initial decrease followed by an increase in average regret. Notably, irrespective of $\beta$ variation, the performance of OTS-RM is always better than UCB-RM. Details are shown in Figure13.
Figure 13: Comparison between UCB-RM and OTS-RM with different choices of $\beta$
Another experiment on testing the sensitivity of OTS-RM is in Figure14. The parameter in OTS-RM is denoted as $M$, which signifies the number of samples taken for optimism in each round. Details about $M$ are outlined in Theorem3.8 and AppendixF.4. Noteworthy from the observations in Figure14 is the minimal impact of varying $M$ across different values $(M=10,20,30)$ on the performance outcome, suggesting adjustments in $M$ do not significantly influence the performance.
Figure 14: Sensitivity test for OTS-RM with different $M$
G.6 Convergence rate related with dimensions
The average regret bounds for IWE-Hedge and the proposed algorithms are $\tilde{\mathcal{O}}(\sqrt{(M+N)/T})$ and $\tilde{\mathcal{O}}(\sqrt{MN/T})$, respectively. Taking OTS-Hedge as an example, for a fixed iteration $T$, IWE-Hedge implies $\log(\text{average regret})\propto(1/2)\log(M+N)$, while OTS-Hedge indicates $\log(\text{average regret})\propto\frac{1}{2}\log(MN)$. When $M=N$, the logarithmic average regret for IWE-Hedge and TS-Hedge should increase with respect to $M$ at the rates of $1/2$ and $1$, respectively. The experimental results shown in Figure 15, where $M$ ranges from $2$ to $100$ with a fixed $T=10^{7}$, match the theoretical predictions almost exactly. These empirical results provide strong support for our regret analysis.
Figure 15: Convergence rate with different dimensions.
G.7 Radar anti-jamming problem
The competition between radar and jammer is an important issue in modern electronic warfare, which can be viewed as a non-cooperative game with two players. This competition occurs at the signal level, where both the radar and jammer can change parameters of their transmitted signals. One representative game form is playing in the frequency domain, as the signal with different carrier frequencies is disjoint.
In our example, we consider the pulse radar and the noise-modulated jammer. The radar transmits pulse signals one by one, with a waiting time interval between consecutive pulses. At the beginning of each pulse, both the radar and jammer transmit their own signals. After a short signal propagation time, each player receives their opponent’s signal and obtains a reward. In our setting, both the radar and jammer have three candidate carrier frequencies, denoted as =$\mathcal{F}=\{f_{1},f_{2},f_{3}\}$. The radar player has three sub-pulses in each radar pulse, and each sub-pulse can choose a different carrier frequency. The action set of the radar is denoted as $\mathcal{A}_{R}=\mathcal{F}\times\mathcal{F}\times\mathcal{F}$, which has a total of $27$ different actions. On the other hand, the jammer player can choose one carrier frequency to transmit the jamming signal and change the carrier frequency for different radar pulses. The action set of the jammer is denoted as $\mathcal{A}_{J}=\mathcal{F}$. After each iteration between the radar and jammer, the radar obtains a signal-inference-and-noise ratio (SINR), which serves as the reward for the radar in that round.
We can observe that both the radar and jammer’s actions are related to the frequency set =$\mathcal{F}=\{f_{1},f_{2},f_{3}\}$. Therefore, the reward in each round can be further defined as a linear function, making the anti-jamming scenario a linear game for the radar.
Definition G.1.
The reward function in the anti-jamming problem is defined as follows:
$\text{SINR}(a,b;\theta)=\phi(a,b)^{T}\theta$
Here, $\phi(a,b)=P_{a}(\theta)/(P_{n_{0}}+P_{b}\mathds{1}(a=b))$ represents a known feature mapping that maps the actions of the radar and the jammer, denoted as $(a,b)$, into the frequency domain related with $\theta$. The parameter $\theta$ corresponds to the radar cross section (RCS) associated with the frequencies in the anti-jamming scenario. $P_{a}(\theta)$ is radar’s received power related with $\theta$, $P_{n_{0}}$ is the noise power and $P_{b}$ is received jammer’s power. The indicator function $\mathds{1}(a=b)$ evaluates to $1$ if the radar and jammer transmit on the same carrier frequency $(a=b)$ and 00 otherwise.
This definition captures the essence of the anti-jamming problem, allowing us to evaluate the performance of different strategies and algorithms based on the signal-to-interference-plus-noise ratio (SINR), and construct the anti-jamming scenario as a linear game.
Figure 16: Illustration of the game between radar and jammer for one pulse.
Table 7: Parameters of FA radar and jammer.
Parameter Value
radar transmitter power $P_{T}$ $30$kW
radar transmit antenna gain $G_{T}$ $32$dB
radar initial frequency $f_{0}$ $3$GHz
bandwidth of each subpulse $B$ $2$MHz
distance between the radar and the jammer $100$km
false alarm rate $p_{f}$ $1\times 10^{-4}$
jammer transmitter power $P_{J}$ $100$W
jammer transmit antenna gain $G_{J}$ $15$dB
Adaptive jammer
In the case of an adaptive jammer, it takes actions according to the radar’s latest 10 pulses. Specifically, it counts the numbers of different carrier frequencies ($f_{1},f_{2},f_{3}\in\mathcal{F}$) appearing in the radar’s last 10 pulses, denoted as $N_{1},N_{2}$, and $N_{3}$. The jammer then takes action $f_{i}$ with a probability proportional to $N_{i}$ for $i=1,2,3$.
G.8 Repeated traffic routing problem
This section considers the traffic routing problem in the transportation literature, which is defined over a directed graph $(\mathcal{V},\mathcal{E})$ ($\mathcal{V}$ and $\mathcal{E}$ are vertices and edges sets respectively) and modeled as a multi-player game. Each node pair (referred as one origin node and one destination node) in the graph is treated as an individual player and every player $i$ seeks to find the ‘best’ route (consists of several edges) to send $U^{i}$ units from its origin node to its destination node. The quality of the chosen route is measured by the travel time, which depends on the total occupancy of the traversed edges. If one edge is occupied by more players, more travel time of this edge is. Specifically, the travel time $t_{e}$ of edge $e\in\mathcal{E}$ is a function of the total units $u$ traversing $e$. One common choice is the Bureau of Public Roads function (Leblanc, 1975)
$t_{e}(u)=c_{e}(1+0.5(\frac{u}{C_{e}})^{4}),$
where $c_{e}$ and $C_{e}$ are free-flow and capacity of edge $e$ respectively. The action of player $i$ is denoted as $a^{i}\in\mathcal{A}^{i}\subset\mathbb{R}^{|\mathcal{E}|}$ and the component corresponding to edge $e$ is denoted as $[a^{i}]_{e}$. If edge $e$ belongs to the route, $[a^{i}]_{e}=U^{i}$ and otherwise $[a^{i}]_{e}=0$. Further, let $a^{-i}\in\mathcal{A}^{-i}$ be the action of other players, the total occupancy of edge $e$ is $[a^{i}]_{e}+[g(a^{-i})]_{e}$, where $g(a^{-i})=\sum_{j\neq i}a^{j}$. This way, the total travel time of a joint action $\bm{a}=(a^{i},a^{-i})$ for player $i$ can be expressed as
$\ell^{i}(a^{i},a^{-i})=\sum_{e\in\mathcal{E}}[a^{i}]_{e}t_{e}([a^{i}]_{e}+[g(a^{-i})]_{e}).$
The reward function of player $i$ is $r^{i}(a^{i},a^{-i})=-\ell^{i}(a^{i},a^{-i})$. Note that the mathematical form of the reward function is unknown to players, only values of $r^{i}(a^{i},a^{-i})$ and actions $(a^{i},a^{-i})$ can be observed by players.
In our experiment, the Sioux-Falls road network data set (Bar-Gera, 2015) is used and we set $c_{e}$ and $C_{e}$ (Bar-Gera, 2015). This network is a directed graph with $24$ nodes and $76$ edges and there are in total $N=528$ players. Each player $i$’s action space $\mathcal{A}^{i}$ is specified by the 5 shortest routes and any route that more than three times longer than the shortest route is further removed from $\mathcal{A}^{i}$. To exploit the correlations among actions $(a^{i},a^{-i})$ in the reward function, the composite kernel proposed in (Sessa et al., 2019) is used. For player $i$, let $a^{i},b^{i}\in\mathcal{A}^{i}$ and $a^{-i},b^{-i}\in\mathcal{A}^{-i}$, then the kernel function associating $(a^{i},a^{-i})$ and $(b^{i},b^{-i})$ is
$K^{i}((a^{i},a^{-i}),(b^{i},b^{-i}))=k_{\rm L}(a^{i},b^{i})k_{\rm P}(a^{i}+g(a^{-i}),b^{i}+g(b^{-i})),$
where $k_{\rm L}(\cdot,\cdot)$ and $k_{\rm P}(\cdot,\cdot)$ are linear and polynomial kernels respectively. The hyperparameters of kernels are set the same as in (Sessa et al., 2019). Details on GP for estimating reward functions can be found in Section 2 of (Sessa et al., 2019) and we do not repeated them here666 The implementation refers the code released by authors of (Sessa et al., 2019) at https://github.com/sessap/stackelucb. Except the regret, the congestion of edge $e$ is also included as a performance metric for a joint action $\bm{a}$, which is computed as $0.15(\sum_{i=1}^{N}[a^{i}]_{e}/C_{e})^{4}$. The averaged congestion is obtained by averaging congestion over all edges.